### v1

# PLAME: Leveraging Pretrained Language Models to Generate Enhanced Protein Multiple Sequence Alignments

Abstract
Protein structure prediction is essential for drug discovery and understanding biological functions. While recent advancements like AlphaFold have achieved remarkable accuracy, most folding models rely heavily on multiple sequence alignments (MSAs) to boost prediction performance. This dependency limits their effectiveness on low-homology proteins and orphan proteins, where MSA information is sparse or unavailable. To address this limitation, we propose PLAME, a novel MSA design model that leverages evolutionary embeddings from pretrained protein language models. Unlike existing methods, PLAME introduces pretrained representations to enhance evolutionary information and employs a conservation-diversity loss to enhance generation quality. Additionally, we propose a novel MSA selection method to effectively screen high-quality MSAs and improve folding performance. We also propose a sequence quality assessment metric that provides an orthogonal perspective to evaluate MSA quality. On the AlphaFold2 benchmark of low-homology and orphan proteins, PLAME achieves state-of-the-art performance in folding enhancement and sequence quality assessment, with consistent improvements demonstrated on AlphaFold3. Ablation studies validate the effectiveness of the MSA selection method, while extensive case studies on various protein types provide insights into the relationship between AlphaFold’s prediction quality and MSA characteristics. Furthermore, we demonstrate that PLAME can serve as an adapter achieving AlphaFold2-level accuracy with the ESMFold’s inference speed.
1 Introduction
Understanding complex and highly variable protein structures is a cornerstone of drug design, enabling breakthroughs in target identification, validation, and drug-target interaction studies baker2001protein ; khoury2014protein ; jumper2021alphafold . Recent advancements, such as AlphaFold, have revolutionized structural biology by achieving near-experimental accuracy for most protein structures and their complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate . However, many of these state-of-the-art models rely heavily on evolutionary information encoded in Multiple Sequence Alignments (MSAs) as input jumper2021alphafold ; abramson2024accurate , making their predictive performance highly dependent on the quality of the MSAs. This dependency poses significant challenges for low-homology protein families and orphan proteins, where evolutionary information is sparse or unavailable kwon2021assessment ; webb2016comparative .
Traditional approaches enhancing low-homology and orphan protein structure prediction include physics-based and template-based methods. Physics-based methods simulate energy functions but are computationally expensive and limited by approximations rohl2004rosetta ; cornell1995amber . Template-based methods rely on sequence homology detection hildebrand2009fast ; finn2011hmmer , making them unsuitable for orphan proteins where evolutionary information is scarce. These limitations have motivated the development of data-driven approaches to improve MSA generation.
Figure 1: Taxonomy of existing MSA designers. Previous methods capture discrete-space distribution of MSAs, including sequence inpainting and prompt-based generation. PLAME serves as an evolutionary-space-inspired de novo generation method independent of sequence prompt.
Recent MSA generation methods can be categorized into sequence space inpainting and prompt-based conditional generation approaches (Figure 1). Sequence inpainting methods like MSA Generator and EvoGen operate within sequence space by completing partial MSAs based on patterns from existing data zhang2023enhancing ; zhang2022few , aiming to reconstruct evolutionary constraints. Prompt-based conditional generation methods, such as MSAGPT and EvoDiff, utilize pretrained models conditioned on MSA prompts to generate new sequences chen2024msagpt ; alamdari2023protein . These AI-based approaches have shown potential in augmenting low-homology MSAs and enhancing structure prediction. As an alternative approach, models like ESMFold construct implicit evolutionary representations directly from single sequences using pretrained protein language models lin2023evolutionary . While this strategy circumvents the need for MSAs, these methods cannot utilize structural templates or enhanced homology information, leading to limited folding accuracy.
Despite recent advancements, MSA generation methods targeting the enhancement of protein folding still face two major challenges. First, these models heavily rely on existing MSA databases, which are biased toward well-represented protein families and provide limited data for low-homology or orphan proteins. This reliance hinders their ability to construct evolutionary spaces that capture the unique constraints of such proteins, reducing their effectiveness in structure prediction. Second, there is a significant gap in understanding how MSA characteristics influence folding model performance. Current methods lack efficient evaluation metrics directly tied to folding performance, creating a disconnect between MSA generation objectives and structural accuracy. Additionally, the response of different protein types to MSA quality varies considerably, yet remains insufficiently studied. While approaches like MSAGPT have attempted to integrate structure prediction feedback through fine-tuning chen2024msagpt , their high computational cost limits scalability. A more comprehensive understanding of the MSA-folding relationship would enable more targeted optimization strategies.
To address these limitations, we propose PLAME, a PLm Aligner for MSA Enhancement designed for low-homology protein structure enhancement. Our key contributions are as follows:
  • Enriched evolutionary spaces: PLAME leverages pretrained protein language models to incorporate evolutionary embeddings as input, providing richer evolutionary information than sparse MSAs for low-homology and orphan proteins wang2024maape ; mcwhite2023leveraging ; hong2024fast . Furthermore, we introduce a novel conservation-diversity loss, which increases model sensitivity to conserved regions while maintaining sequence diversity, enabling the generation of biologically plausible MSAs optimized for high-quality alignments.
  • Comprehensive MSA quality exploration: We propose a novel sequence quality assessment metric that provides a new perspective for evaluating MSA quality and its relationship to folding performance. Based on these insights, we develop an MSA selection method called High-Fidelity Appropriate Diversity (HiFiAD) that improves folding accuracy by effectively screening MSAs that balance sequence fidelity and diversity. This approach offers both analytical understanding and practical improvement for protein structure prediction.
  • Extensive experimental validation: Our experimental results demonstrate the substantial enhancement on challenging low-homology and orphan protein datasets when evaluated with both AlphaFold2 and AlphaFold3 folding frameworks. Ablation studies confirm the effectiveness of the conservation-diversity loss and HiFiAD selection method in improving both MSA quality and folding performance. Additionally, we conduct comprehensive experiments on multiple protein systems, including general and de novo proteins, providing deeper insights into the relationship between MSA characteristics and folding performance.
Figure 2: Overview of PLAME framework. PLAME captures ESM-2 evolutionary representations, generating MSAs for augmenting the original MSAs. The augmented MSAs serve as the homology template for folding softwares for folding enhancement. In each block of the T5-architecture, additional row-attention and col-attention are applied to capture co-evolutionary information.
2 Related Works
Protein Structure Prediction
Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches. Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding cornell1995amber ; brooks2009charmm . While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results karplus2002molecular ; freddolino2010challenges ; pande2010everything . Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures rohl2004rosetta ; hildebrand2009fast . These methods perform well with suitable templates but struggle with orphan proteins and low-homology families webb2016comparative ; baker2001protein . Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability jumper2021alphafold ; abramson2024accurate ; wu2022high . Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
AlphaFold-based Enhancement
Building on AlphaFold’s success, researchers have developed methods to refine specific modules, aiming to improve accuracy or efficiency. These advancements can be grouped into three main categories. The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy. However, these methods often slow down inference despite their modest performance gains johnson2010hidden ; steinegger2017mmseqs2 ; zheng2024improving ; lee2024petabase . The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency. However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy lin2023evolutionary ; mirdita2022colabfold . The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families. While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability alamdari2023protein ; zhang2022few ; zhang2023enhancing ; chen2024msagpt .
3 Method
3.1 Problem Formulation
Protein structure prediction relies heavily on high-quality MSAs to provide evolutionary information, but the accuracy of folding software $\mathcal{F}_{\omega}$ significantly drops when MSAs are sparse or insufficient. Given proteins $\mathbf{P}=\{\mathbf{s},\mathbf{x},\mathbf{M}\}$, where $\mathbf{s}\in\mathcal{S}$ are query sequences, $\mathbf{x}\in\mathcal{X}$ are 3D structures, and $\mathbf{M}=\{\mathbf{m}_{1},\mathbf{m}_{2},\dots,\mathbf{m}_{N}\}\in\mathcal{M}$ are MSAs with each $\mathbf{m}_{i}$ as an aligned homologous sequence. The goal of MSA design models $\mathbf{p}_{\theta}:\mathcal{M}\rightarrow\mathcal{M}$ is designing augmented MSAs $\mathbf{M}_{\text{aug}}$ that enhances evolutionary information to obtain more accurate structures $\mathbf{x}^{\prime}$ using folding software $\mathcal{F}_{\omega}$.
$\mathbf{M}^{\prime}=\mathbf{p}_{\theta}(\mathbf{M}),\quad\mathbf{x}^{\prime}=\mathcal{F}_{\omega}(\mathbf{s},\mathbf{M}_{\text{aug}})$ (1)
where the augmented MSAs are composed of original MSAs $\mathbf{M}$ and generated MSAs $\mathbf{M}^{\prime}$, denoted as $\mathbf{M}_{\text{aug}}=\{\mathbf{M},\mathbf{M}^{\prime}\}$. The quality of the enhanced structures is evaluated using several metrics, including RMSD, TM-score, and pLDDT.
The key to high-fidelity MSA generation lies in constructing an informative evolutionary distribution $\mathbf{z}_{\text{evo}}$, which serves as the foundation for generating augmented MSAs $\mathbf{M}_{\text{aug}}$. Current methods utilize deep neural networks $\mathbf{f}_{\theta}$ to learn hidden evolutionary distributions directly from existing MSAs.
$\mathbf{z}_{\text{evo}}=\mathbf{f}_{\theta}(\mathbf{M})$ (2)
However, relying solely on sequence-level information from MSAs fails to capture the complete evolutionary landscape, particularly when MSA coverage is sparse or incomplete. To overcome this limitation, we propose an evolutionary space based on evolutionary embeddings derived from pretrained protein language models (PLMs) $\mathbf{g}_{\phi}$.
$\mathbf{z}_{\text{evo}}=\mathbf{f}_{\theta}(\mathbf{g}_{\phi}(\mathbf{s}))$ (3)
3.2 Model Architecture
PLAME employs an encoder-decoder transformer architecture similar to MSA Transformer rao2021msa , with adjustments to the T5 block structure vaswani2017attention . The encoder and decoder incorporate additional row-wise and column-wise attention mechanisms, designed to better capture evolutionary patterns in MSA data (detailed in Fig 2). Additional mechanisms are introduced as follows.
Row Attention
The row attention mechanism models inter-sequence dependencies within the evolutionary space, summarizing sequence relationships across the depth of the input representation. Given an input tensor $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{L\times N\times D\times h}$, where $L$ is the sequence length, $N$ is the number of MSAs, $D$ is the depth, and $h$ is the hidden dimension, a global representation $\mathbf{H}r$ is computed by averaging the hidden states along the depth dimension:
$\mathbf{H}_{r}=\frac{1}{D}\sum_{d=1}^{D}\mathbf{H}_{\text{enc}}^{d},\quad\mathbf{H}_{r}\in\mathbb{R}^{L\times N\times h}.$ (4)
Here, $\mathbf{H}_{r}$ encodes the evolutionary space, facilitating cross-attention during decoding. The cross-row attention (Row-Attn) is defined as:
[…]
where $\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r}$ denote the Query, Key, and Value matrices, respectively.
Column Attention
Column attention captures conservation patterns across columns in the MSA, focusing on evolutionary signals at specific positions. To implement this mechanism, the input representation matrix $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{D\times N\times L\times h}$ is first transposed to $\mathbf{X}_{\text{dec}}^{\top}\in\mathbb{R}^{L\times N\times D\times h}$. The self and cross-column attention mechanisms operate on this transformed representation.
For cross-column attention, the query, key, and value matrices $\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}}$ and their corresponding projection weights $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are defined as:
$\mathbf{Q}_{\text{c}}=\mathbf{X}_{\text{dec}}^{\top}\mathbf{W}_{q},\quad\mathbf{K}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{k},\quad\mathbf{V}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{v}.$ (6)
The column attention is computed as:
[…]
Generation & Inference
In our implementation, ESM-2 ($\mathbf{g}_{\phi}$) encodes the query sequence $\mathbf{s}$ into high-dimensional evolutionary embeddings, denoted as $\mathbf{H}_{\text{input}}$. These embeddings are processed by the encoder through $N$ layers of modified T5 blocks, iteratively refining contextualized representations:
$\mathbf{H}_{\text{Enc}}^{(l)}=\mathbf{Enc}^{(l)}(\mathbf{H}^{(l-1)}),\quad l=1,\dots,N,\quad\mathbf{H}^{(0)}=\mathbf{H}_{r}.$ (8)
The final encoder output, $\mathbf{H}^{(N)}$, captures comprehensive sequence context and is passed to the decoder for autoregressive generation. The decoder generates tokens sequentially, conditioned on the encoder output $\mathbf{H}_{\text{Enc}}^{(N)}$ and previously generated tokens:
$\mathbf{y}_{t}=\mathbf{Dec}(\mathbf{y}_{<t},\mathbf{H}_{\text{Enc}}^{(N)}).$ (9)
The T5 blocks with row and column attention enable the decoder to compute output embeddings, which are then passed through a softmax layer to produce token probabilities for the next position.
3.3 Conservation-Diversity Training Loss
We introduce a position-aware causal inference approach for diverse MSA generation, integrating a PSSM-Weighted Cross-Entropy Loss and a Diversity Regularization Loss to balance focus on conserved regions with sampling diversity.
PSSM-Weighted Cross Entropy (PCE) Loss
The PSSM-Weighted Cross Entropy Loss emphasizes accurate predictions in conserved regions of the MSA, which are critical for maintaining protein structure and function. For a single sequence, it is defined as:
$\mathcal{L}_{\text{seq}}=-\sum_{l=1}^{L}w_{l}\cdot\log p(y_{l}\mid y_{<l}),$ (10)
where $L$ is the sequence length, $y_{l}$ is the true label at position $l$, and $p(y_{l}\mid y_{<l})$ is the predicted probability of $y_{l}$.
The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position. These weights are normalized to the range $[1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation. Specifically,
$w_{l}=1+\delta\cdot\frac{\text{freq}_{l}-\text{min}(\text{freq})}{\text{max}(\text{freq})-\text{min}(\text{freq})}.$ (11)
During model training, we apply $\delta=0.5$, assigning higher weights to conserved positions and lower weights to less conserved ones. For a batch of $N$ sequences, the PCE loss averages over all sequences and positions:
$\mathcal{L}_{\text{PCE}}=-\frac{1}{N}\sum_{j=1}^{N}\sum_{l=1}^{L_{j}}w_{l}^{(j)}\cdot\log p(y_{l}^{(j)}\mid y_{<l}^{(j)}),$ (12)
where $L_{j}$ is the length of the $j$-th sequence, and $w_{l}^{(j)}$ is the weight for position $l$ in sequence $j$. This loss emphasizes conserved regions while allowing flexibility in less conserved areas.
Diversity Regularization (DIRE) Loss
The Diversity Regularization Loss promotes biological diversity in MSAs by maximizing the entropy of predicted amino acid distributions. For a single sequence, the entropy at position $l$ is calculated as:
$H_{l}=-\sum_{a\in\mathcal{A}}p(a\mid y_{<l})\log p(a\mid y_{<l}),$ (13)
where $p(a\mid y_{<l})$ is the predicted probability of amino acid $a$ at position $l$, and $\mathcal{A}$ is the set of all amino acids. To encourage diversity across a batch of $N$ sequences, we compute the average entropy over all positions and sequences:
$\mathcal{L}_{\text{diversity}}=-\frac{1}{N}\sum_{j=1}^{N}\frac{1}{L_{j}}\sum_{l=1}^{L_{j}}H_{l}^{(j)},$ (14)
where $L_{j}$ is the length of sequence $j$, and $H_{l}^{(j)}$ is the entropy at position $l$ in sequence $j$. This loss encourages the model to capture the natural amino acid diversity in homologous sequences.
Combined Loss Function
The combined loss function balances conservation and diversity:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{\text{PCE}}+(1-\alpha)\cdot\mathcal{L}_{\text{DIRE}},$ (15)
where $\alpha\in[0,1]$ controls the trade-off between the two terms. We set $\alpha=0.9$ in our experiments, prioritizing conservation while maintaining sufficient diversity. This design ensures the generated MSAs accurately capture conserved regions and retain natural sequence variability. Our theoretical analysis in Appendix A demonstrates that PCE Loss enhances the model’s understanding of evolutionary information (MSA profile), while DIRE Loss functions as a regularizer to prevent neglect of variable regions.
3.4 MSA Selection Method – HiFiAD
Existing MSA generation models struggle in ensuring the quality of generated MSAs. Different methods capture MSA distributions with varying effectiveness, and excessive concatenation of virtual MSAs can degrade the quality of the original MSA. This, in turn, restricts the ability of protein structure prediction software to interpret evolutionary distributions. While MSAGPT has explored MSA selection in its ablation studies, existing approaches lack a clear definition of MSA quality and fail to propose systematic selection methods based on sequence-level quality metrics chen2024msagpt .
During MSA augmentation, virtual MSAs inevitably introduce noise, limiting the performance of folding software. A common issue with current models is the generation of over-conserved sequences that closely resemble the query sequence. When these sequences are over-concatenated, they distort the evolutionary distribution of the original MSA. This issue was highlighted in MSAGPT’s studies, which showed that performance plateaued once the number of generated MSAs exceeded 32. To address this, balancing recovery rate distributions by selecting sequences with both high and low recovery rates has proven more effective than focusing exclusively on one or the other.
To address these issues, we propose HiFiAD to balance fidelity and diversity for MSA selection while maintaining efficiency. For orphan proteins without original MSAs (termed zero-shot cases), we select the top-$k$ sequences with the highest $S_{\text{BLOSUM}}$, along with sequences from the top and bottom $k/2$ of the recovery rate distribution. The BLOSUM substitution score is calculated as:
$S_{\text{BLOSUM}}(m_{i},s)=\sum_{j=1}^{L}B(s_{j},m_{ij}),\quad\forall m_{i}\in M,$ (16)
where $B$ is the BLOSUM matrix (BLOSUM62 in our case). For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution. Specifically, we cap the number of MSA sequences $N_{\text{aug}_{\text{max}}}$ at $\max(16,2N)$. If fewer than $N_{\text{aug}_{\text{max}}}$ virtual MSAs are available, we retain all generated MSAs to ensure sufficient diversity for downstream analysis.
Table 1: Performance metrics across different modes and models. The best results in each folding mode are highlighted in bold. Zero and Few indicate zero-shot (proteins without MSAs) and few-shot cases (proteins with existing MSAs), respectively.
pLDDT GDT TMscore RMSD LDDT pTM
Zero Few Zero Few Zero Few Zero Few Zero Few Zero Few
ESMFold 66.26 62.62 0.6 0.53 0.6 0.57 9.58 12.04 0.62 0.59 / /
Mode1
AF2 MSA 60.07 62.14 0.50 0.52 0.50 0.57 12.34 12.16 0.54 0.58 0.44 0.49
EvoDiff 58.68 61.83 0.46 0.50 0.46 0.54 13.81 12.95 0.50 0.56 0.40 0.48
MSAGPT 59.81 61.18 0.48 0.51 0.48 0.56 12.62 12.35 0.53 0.57 0.43 0.48
PLAME 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
Mode2
AF2 MSA 66.56 66.32 0.51 0.55 0.52 0.60 12.06 11.84 0.55 0.61 / /
EvoDiff 61.98 65.83 0.48 0.53 0.48 0.58 14.23 11.82 0.52 0.59 / /
MSAGPT 64.88 65.96 0.51 0.56 0.51 0.60 12.60 11.90 0.55 0.61 / /
PLAME 67.77 67.48 0.53 0.55 0.54 0.60 12.62 11.90 0.57 0.61 / /
Mode3
AF2 MSA 70.31 69.61 0.57 0.60 0.57 0.64 10.53 10.24 0.60 0.65 / /
EvoDiff 64.39 68.54 0.51 0.57 0.51 0.61 13.20 10.81 0.54 0.62 / /
MSAGPT 68.39 69.30 0.57 0.60 0.56 0.64 11.05 10.40 0.59 0.64 / /
PLAME 71.50 70.48 0.58 0.59 0.58 0.64 11.41 10.62 0.60 0.64 / /
4 Experiment
Baselines
To evaluate PLAME’s capability in generating high-fidelity and diverse MSAs, we compared it with several state-of-the-art AI-based MSA generation methods and AlphaFold2’s MSA pipeline jumper2021alphafold . The baselines include AF2 MSA, and open-source methods including EvoDiff and MSAGPT chen2024msagpt ; alamdari2023protein . Additionally, to assess the potential advantages of leveraging AlphaFold2’s knowledge base in protein structure prediction, we included ESMFold lin2023evolutionary , an MSA-free folding software, as reference.
Evaluation Metric
We apply the following metrics for structure assessment: 1) local metrics including pLDDT and LDDT; 2) global metrics including GDT, TM-Score, pTM, and RMSD. To comprehensively evaluate the effectiveness of augmented MSAs, we tested three AF2 folding mode, progressing from computationally efficient, basic settings to more comprehensive configurations: 1) Mode1: Folding using pTM-3 model without using templates; 2) Mode2: Folding using the default 5 models without using templates; 3) Mode3: Folding using the default 5 models with templates. To evaluate the quality of generated multiple sequence alignments (MSAs), we employ four sequence-based metrics to quantify alignment fidelity and diversity: 1) Conservation Score measures the degree of residue conservation at each alignment position. It is defined as $C_{i}=\frac{\text{Freq}_{\text{max}}(i)}{N}$, where $\text{Freq}_{\text{max}}(i)$ is the frequency of the most common residue at position $i$, and $N$ is the total number of sequences. A high score indicates strong conservation, often reflecting functional or evolutionary constraints, while a low score suggests positional variability. 2) Gap Proportion quantifies the fraction of gaps at each alignment position, calculated as $G_{i}=\frac{G(i)}{N}$, where $G(i)$ is the number of gaps at position $i$. Lower values indicate better alignment quality with fewer missing data or alignment errors. 3) Substitution Compatibility assesses the evolutionary plausibility of aligned residues using BLOSUM62 substitution scores $S_{\text{BLOSUM}}$ (Eq. 16). Higher scores indicate evolutionary compatibility and alignment biological relevance. 4) Alignment Entropy captures residue diversity at each position using Shannon entropy:
$H_{i}=-\sum_{r\in\{R_{i}\}}p(r)\log_{2}p(r),$
where $\{R_{i}\}$ is the set of unique residues at position $i$, and $p(r)=\frac{\text{count}(r)}{N}$ is the relative frequency of residue $r$. Higher entropy signifies greater positional diversity, indicating weaker functional constraints, whereas lower entropy denotes stronger conservation, often associated with functional or evolutionary significance.
Datasets
For the training dataset, we use the PDB and UniClust30 subsets from the open-source OpenProteinSet as our data source ahdritz2024openproteinset . The pre-searched MSAs from OpenFold training are also included. We only retain data with at least 64 MSA sequences. To avoid overlap with the test cases, we removed sequences with over 90% similarity using MMSeqs based on UniClust30 clustering results mirdita2017uniclust ; steinegger2017mmseqs2 . This process yields an initial dataset of 293,979 samples, which are split into training and validation sets with a 90:10 ratio. For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein . Any $>90\%$ redundancy between the test cases and training dataset is eliminated.
4.1 PLAME as High-quality MSA Designer
General folding performance
We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics. Results are shown in Table 1. We analyze the quality of MSAs generated by different methods and their performance in various scenarios. 1) Performance Superiority: PLAME achieved state-of-the-art performance on most metrics across all modes. This demonstrates that PLAME-generated MSAs effectively enhance AF2 folding, particularly in scenarios involving low-homology or orphan proteins, where significant improvements in prediction accuracy were observed. In contrast, EvoDiff and MSAGPT struggled to produce high-quality MSAs under these conditions. By leveraging the evolutionary latent space provided by ESM-2, PLAME generated biologically meaningful virtual MSAs, outperforming baselines on key metrics such as pLDDT and RMSD. 2) Diminishing Returns: As more advanced folding configurations were employed (e.g., with more powerful AF2 models or the inclusion of templates), the benefits of MSA augmentation gradually decreased. In Mode3, where templates were used, certain metrics exhibit performance decrease. This is likely because template information already captures much of the necessary sequence and evolutionary information, reducing the marginal utility of additional MSA augmentation. Furthermore, the noise introduced by low-quality MSAs can be amplified in stronger baselines, further impacting performance. 3) Metric Discrepancies: The enhancement effects across different metrics were not entirely consistent. For instance, the trends for pLDDT and RMSD diverged. While pLDDT, as a confidence metric, reflects the model’s confidence in its predictions, RMSD measures the global deviation of the predicted structure from the ground truth. MSA augmentation improved local prediction quality (leading to higher pLDDT), but had limited impact on reducing global structural deviation (resulting in minimal RMSD improvement). Nevertheless, pLDDT alone is sufficient to demonstrate the high quality of the generated MSAs. 4) Noise Issues: EvoDiff and MSAGPT produced MSAs that consistently underperformed compared to the original AF2 MSAs across all modes. This was particularly evident in low-homology protein scenarios, where the generated MSAs failed to accurately capture evolutionary distributions and often included irrelevant sequences. When concatenated with the original AF2 MSAs, these low-quality sequences introduced additional noise, disrupting the model’s inference. This aligns with the hypothesis presented in Section 3.4, where low-quality MSAs were predicted to introduce noise, ultimately impairing folding performance. 5) Few/Zero-shot Consistency: In Few-shot scenarios, EvoDiff and MSAGPT performed relatively better due to the presence of initial MSAs that provided additional homology information to guide the generation process. In contrast, PLAME does not rely on initial MSA searches and directly generates high-quality MSAs through its evolutionary latent space. This enables PLAME to excel in Zero-shot scenarios while maintaining consistent performance in Few-shot cases. This advantage significantly reduces runtime costs while ensuring robust enhancements across all scenarios.
Sequence quality assessment
Figure 3: Comparison of sequence-based metrics for AF2 MSAs and MSAs generated by EvoDiff, MSAGPT, and PLAME.
To evaluate generated MSA quality, we designed a set of metrics focusing on MSA fidelity and diversity. These metrics aim to investigate which factors correlate most strongly with protein folding performance, especially given the lack of standardized criteria for such evaluations. Our study provides an initial exploration into this problem. Figure 3 illustrates the evaluation results.
PLAME-generated MSAs outperform other methods on fidelity metrics, with distributions closest to AF2 MSAs in Conservation Score, Gap Proportion, and Substitution Compatibility. Specifically, PLAME achieves higher Conservation Scores and Substitution Compatibility, reflecting its ability to better capture evolutionary and functional constraints. Moreover, its lower Gap Proportion indicates higher alignment completeness, which can be attributed to the latent evolutionary space providing richer homology information. These results highlight PLAME’s superior alignment with the evolutionary constraints of the target protein.
We analyzed Alignment Entropy as a measure of diversity. While greater diversity is generally expected to enhance homologous information, our results show that PLAME’s diversity levels are closer to those of AF2 MSAs, rather than exceeding methods like EvoDiff. This supports our assumption in 3.4 that excessive diversity can introduce noise and diminish the information enrichment effect of generated MSAs. Therefore, achieving a balance between fidelity and moderate diversity is crucial for MSA generation models that struggle to capture high-quality distributions.
4.2 PLAME as MSA-free ESMFold Adapter to AlphaFold2
The comparative analysis between PLAME and ESM-Fold demonstrates PLAME’s growing advantage as the folding modes become more advanced. In Mode1 and Mode2, PLAME outperforms ESMFold in key metrics such as pLDDT and RMSD, with obvious gains in pLDDT. In Mode3, the introduction of structural templates amplifies PLAME’s advantage, achieving better results across all metrics. For instance, pLDDT of PLAME increases from 66.54 to 71.50, surpassing ESMFold’s 66.26. These results validate PLAME’s ability to merge ESMFold’s computational efficiency with AF2’s performance by generating high-quality virtual MSAs. Acting as an adaptive network, PLAME transforms ESMFold’s single-sequence inference into AF2’s MSA-based folding, integrating evolutionary information to enhance prediction accuracy. Thus, PLAME overcomes ESMFold’s limitations by introducing virtual MSA generation, achieving significant improvements in Mode3 and providing an efficient and accurate solution for protein structure prediction.
Table 2: Ablation study over different MSA selection approaches.
pLDDT GDT TMscore RMSD LDDT pTM
Zero Few Zero Few Zero Few Zero Few Zero Few Zero Few
Random-16 63.61 62.63 0.52 0.51 0.52 0.56 12.01 12.67 0.55 0.58 0.46 0.49
Blosum-8 61.04 62.71 0.5 0.52 0.51 0.57 12.53 12.69 0.55 0.58 0.45 0.50
Blosum-32 62.97 62.40 0.50 0.50 0.51 0.55 12.28 12.84 0.55 0.57 0.45 0.48
Top-Rec-16 62.04 62.93 0.51 0.51 0.51 0.55 12.15 12.48 0.55 0.57 0.45 0.49
Top-down-Rec-16 63.43 63.10 0.52 0.52 0.51 0.57 11.97 12.15 0.55 0.58 0.46 0.49
HiFiAD 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
4.3 Ablation on Selection Methods
To evaluate the impact of MSA quality (3.4), we perform an ablation study using different MSA selection strategies. The results are summarized in Table 2. The selection methods are based on sequence similarity metrics (Rec) and BLOSUM substitution scores (BLOSUM). Additionally, Top and Down refer to the highest- and lowest-scoring cases. The experiments lead to the following key observations. 1) Effective HiFiAD filtering: HiFiAD achieves superior performance across all metrics compared to other filtering methods, demonstrating that the combination of high fidelity and balanced diversity is essential. 2) High-fidelity: Compared to the Top-down-Rec method and random selection, HiFiAD identifies more high-fidelity cases, validating the importance of leveraging in-distribution MSAs. 3) Diverse samples: Compared to BLOSUM-based methods and Top-Rec, HiFiAD selects more diverse samples. This balanced diversity prevents the conservation distribution from being overly deterministic, which could potentially correct the augmented MSA distribution. Among sequence-metric-based filtering methods, HiFiAD dynamically adjusts to Few- and Zero-shot scenarios, as well as varying MSA quality levels, making it a simple yet effective approach for MSA selection. More ablation results are shown in Table 5 in Appendix.
5 Conclusion
In this work, we introduced PLAME, the first model to leverage evolutionary embeddings for MSA sequence generation. Our approach bridges the gap between single-sequence inference and MSA-based methods, effectively enhancing protein folding performance. Our extensive evaluations demonstrate that PLAME-generated MSAs surpass existing methods in both conservation and diversity metrics, leading to significant improvements in structure prediction accuracy across various protein families. By functioning as both an MSA augmenter and an efficient AlphaFold adapter without the need for time-intensive MSA searches, PLAME establishes itself as a fast, accurate, and scalable solution for protein structure prediction. Furthermore, our proposed quality metrics and comprehensive experiments provide new insights into the relationship between MSA characteristics and folding performance, advancing our understanding of how sequence information translates to structural accuracy.
Appendix A Proof of Theorem
We provide additional statements to demonstrate the superiority of the Conservation-Diversity Training Loss. Firstly, we demonstrate that the PCE Loss as a conservation-aware weighted loss by position in the perspective of MSA profiles.
Lemma 1.
[…]
Proof.
For a sufficiently large set of $N$ homologous sequences sampled from $P$, the expected cross-entropy loss is
$\mathbb{E}[\mathcal{L}_{\text{CE}}]=-\sum_{l=1}^{L}\sum_{a\in\mathcal{A}}P(l,a)\,\log Q_{\theta}(l,a).$ (17)
Re-expressing each column term as $-\sum_{a}P\log Q=H\!\bigl{(}P(l,\!\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}$, we obtain
$\mathbb{E}[\mathcal{L}_{\text{CE}}]=\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}+\sum_{l=1}^{L}H\!\bigl{(}P(l,\!\cdot)\bigr{)}.$ (18)
For the PCE loss,
$\mathbb{E}[\mathcal{L}_{\text{PCE}}]=-\sum_{l=1}^{L}w_{l}\sum_{a\in\mathcal{A}}P(l,a)\,\log Q_{\theta}(l,a),$ (19)
which can analogously be rewritten as the position-wise weighted KL
$\mathbb{E}[\mathcal{L}_{\text{PCE}}]=\sum_{l=1}^{L}w_{l}\,\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}+\sum_{l=1}^{L}w_{l}\,H\!\bigl{(}P(l,\!\cdot)\bigr{)}.$ (20)
Let $\theta$ denote the model parameters. The gradient of the CE loss for column $l$ is
$\frac{\partial\mathcal{L}_{\text{PCE},l}}{\partial\theta}=-\sum_{a\in\mathcal{A}}P(l,a)\,\frac{1}{Q_{\theta}(l,a)}\frac{\partial Q_{\theta}(l,a)}{\partial\theta}.$ (21)
For PCE the gradient is simply scaled by $w_{l}$:
$\frac{\partial\mathcal{L}_{\text{PCE},l}}{\partial\theta}=-w_{l}\sum_{a\in\mathcal{A}}P(l,a)\,\frac{1}{Q_{\theta}(l,a)}\frac{\partial Q_{\theta}(l,a)}{\partial\theta}=w_{l}\,\frac{\partial\mathcal{L}_{\text{CE},l}}{\partial\theta}.$ (22)
Consequently, in highly conserved columns the gradient magnitude is amplified by $1+\delta$, whereas in variable columns ($w_{l}\approx 1-\delta$) it is attenuated, focusing optimization effort on conserved regions. ∎
Based on the understanding of the PCE Loss, we then demonstrate that PCE Loss is expected to capture evolutionary information (MSA profile) with less error–measured by KL-Divergence.
Theorem 1.
Let $P(l,a)$ be the true amino–acid distribution in column $l\;(l=1,\dots,L)$ of an MSA and let $Q_{\theta}(l,a)$ be the distribution produced by a parametrised generative model $Q_{\theta}$. Denote the column–wise Kullback–Leibler divergence by
$\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\,\|\,Q_{\theta}(l,\cdot)\bigr{)}=\sum_{a\in\mathcal{A}}P(l,a)\,\log\frac{P(l,a)}{Q_{\theta}(l,a)}.$ (23)
Let
$\theta^{\star}_{\mathrm{CE}}=\arg\min_{\theta}\mathcal{L}_{\text{CE}}(\theta),\qquad\theta^{\star}_{\mathrm{PCE}}=\arg\min_{\theta}\mathcal{L}_{\text{PCE}}(\theta).$ (24)
Define the average profile KL divergence
$D_{\text{KL}}^{\text{avg}}(\theta):=\frac{1}{L}\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\,\|\,Q_{\theta}(l,\cdot)\bigr{)}.$ (25)
Under the assumption that both optimization problems are solved to global optimality, the model trained with PCE Loss captures the MSA profile with less divergence $D_{\text{KL}}^{\text{avg}}$:
$D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{PCE}})\;\leq\;D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{CE}})$ (26)
Proof.
Rewrite two losses in the form of KL-Divergence $\sum_{a}P\log Q=H\bigl{(}P(l,\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)}$, we have:
$\mathcal{L}_{\text{CE}}(\theta)$ $=C_{0}+\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)},$ (27)
$\mathcal{L}_{\text{PCE}}(\theta)$ $=C_{w}+\sum_{l=1}^{L}w_{l}\,\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)},$
where $C_{0}=\sum_{l}H(P(l,\cdot))$ and $C_{w}=\sum_{l}w_{l}\,H(P(l,\cdot))$ are constants independent of $\theta$. Hence minimizing $\mathcal{L}_{\text{PCE}}$ is equivalent to minimizing the weighted KL
$D_{w}(\theta):=\sum_{l=1}^{L}w_{l}\,\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)},\qquad\theta^{\star}_{\mathrm{PCE}}=\arg\min_{\theta}D_{w}(\theta).$ (28)
Then, since every $w_{l}$ is bounded, we can establish the relations:
$(1-\delta)\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)}\;\leq\;D_{w}(\theta)\;\leq\;(1+\delta)\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\cdot)\|Q_{\theta}(l,\cdot)\bigr{)}.$ (29)
Dividing by $L$ gives:
$(1-\delta)\,D_{\text{KL}}^{\text{avg}}(\theta)\;\leq\;\frac{D_{w}(\theta)}{L}\;\leq\;(1+\delta)\,D_{\text{KL}}^{\text{avg}}(\theta).$ ($\ast$)
Based on the fact that $\theta^{\star}_{\mathrm{PCE}}$ minimizes $D_{w}$, denote $\Delta_{w}:=D_{w}(\theta^{\star}_{\mathrm{CE}})-D_{w}(\theta^{\star}_{\mathrm{PCE}})\geq 0$. By applying ()($\ast$)( ∗ ) to both optimal parameters and subtracting, we obtain:
$(1-\delta)\Bigl{[}D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{CE}})-D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{PCE}})\Bigr{]}\;\leq\;\frac{\Delta_{w}}{L}.$ (30)
Since $\Delta_{w}\geq 0$ and $01-\delta>$ 0; it is strictly positive whenever $\Delta_{w}>0$, Therefore,
$D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{PCE}})\;\leq\;D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{CE}}),$ (31)
which completes the proof. ∎
A natural challenge emerges when applying the PCE Loss—the model tends to accurately capture the distribution of conserved regions while neglecting the distribution of variable regions. To address this issue, we demonstrate that the DIRE Loss effectively enhance the modeling in the variable regions.
Theorem 2.
For $l=1,\dots,L$ let $P(l,a)$ denote the empirical amino-acid distribution and $Q_{\theta}(l,a)$ any model. When each amnio acid site is optimized independently, the minimizer is
$Q_{\alpha}^{\star}(l,a)=\frac{P(l,a)^{\tau_{l}}}{\displaystyle\sum_{b\in\mathcal{A}}P(l,b)^{\tau_{l}}},\qquad\tau_{l}=\frac{\alpha w_{l}}{\alpha w_{l}+(1-\alpha)}\in(0,1).$ (32)
Moreover,
$H\!\bigl{(}P(l,\cdot)\bigr{)}\leq H\!\bigl{(}Q_{\alpha}^{\star}(l,\cdot)\bigr{)}\leq\log|\mathcal{A}|,$ (33)
with the entropy increase largest when $w_{l}$ is small (variable regions). Thus $\mathcal{L}_{\text{DIRE}}$ counter-acts the entropy suppression of $\mathcal{L}_{\text{PCE}}$ and serves as a principled regularizer on variable regions.
Proof.
Since the combined loss $\mathcal{L}_{\alpha}$ sums over amino acid positions, we may analyze a single site independently, denoting $P(a)=P(l,a)$, $Q(a)=Q(l,a)$ and $w=w_{l}$. For each site we minimize, we have
$F(Q)=\alpha w\sum_{a}P(a)\log\frac{P(a)}{Q(a)}+(1-\alpha)\sum_{a}Q(a)\log Q(a),$ (34)
subject to the normalization constraint $\sum_{a}Q(a)=1$.
Introducing a Lagrange multiplier $\lambda$ and setting the derivative with respect to $Q(a)$ to zero yields
[…]
Solving this equation reveals a "temperature-like" solution based on $\tau$:
$Q(a)\;\propto\;P(a)^{\tau},\qquad\tau\;=\;\frac{\alpha w}{\alpha w+(1-\alpha)}\in(0,1),$ (36)
which is exactly the optima $Q_{\alpha}^{\star}(l,\cdot)$ mentioned earlier.
Since $10<\tau<$ 1, this transformation always increases entropy unless $P$ is already uniform:
$H\!\bigl{(}P(l,\cdot)\bigr{)}\;\leq\;H\!\bigl{(}Q_{\alpha}^{\star}(l,\cdot)\bigr{)}\;\leq\;\log|\mathcal{A}|.$ (37)
The entropy gain is larger when $w$ is small (in the variable regions). Consequently, the $(1-\alpha),\mathcal{L}_{\text{DIRE}}$ term counteracts the over-confidence induced by $\mathcal{L}_{\text{PCE}}$ in variable regions, serving as an adaptive entropy-based regularizer. ∎
Appendix B Training and Sampling Details
Training Details
We trained our model based on a Transformer T5 architecture, incorporating axial attention and task-specific modifications to enhance performance. The model consists of 12 encoder layers and 12 decoder layers, with a hidden size of 1024, 12 attention heads, and a feedforward dimension of 2048. The feedforward projection employs a gated-GELU activation function. During training, we employed the AdamW optimizer with a learning rate of 5e-5, a weight decay of 1e-5, and a polynomial decay scheduler with a 1% warmup ratio. Training was conducted on four NVIDIA A40 GPUs for up to 200,000 steps, with a batch size of 4 per device for both training and evaluation.
Sampling details
The sampling process was configured with the following parameters: we generate 16 MSAs for 4 trials per generation. The sampling used a repetition penalty of 1.0, a temperature of 1.0, and top-p sampling with a threshold of 0.95. Beam search was performed with 4 beams and 1 beam group. Sampling was executed on an A40 GPU.
Appendix C Extensive Experimental Results
C.1 Case Study on Successful Designs
To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structrure visualization in Appendix F. We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.
Among these cases, we can generally observe that most generated MSA sequences maintain high similarity with the query sequence. Furthermore, the generated MSAs provide good enhancement at the originally conserved sites. This indicates that protein language models can still retain some evolutionary information even for proteins with low homology, although the diversity they can provide is more limited due to homology constraints.
Additionally, we identified several patterns in the sampled MSAs that clearly deviate from the original distribution, such as consecutive gaps (in 8ehb_F), repeated HHHHHH sequences (in 8okw_B), and repeated SSSSSSSSS (in 7xrl_A). We believe these erroneous generations are related to the autoregressive generation method, where the model tends to produce excessive hallucinations after getting trapped in incorrect local probability distributions. We also observed that these failure patterns occur more frequently in longer sequences, possibly due to insufficient training on cases with greater length. These represent an area requiring further improvement.
C.2 AlphaFold3 Folding Experiment
To further validate PLAME is an universal MSA generation framework for different folding software, we conduct extensive experiment on AlphaFold3 [5]. We use the same generated MSAs filtered by HiFiAD, augmenting the original MSAs for folding enhancement.
pLDDT GDT TMscore RMSD LDDT pTM
AF2 MSA 68.872 0.578 0.596 10.818 0.617 0.529
PLAME 70.887 0.578 0.595 10.740 0.623 0.539
Table 3: Comparison of folding enhancement based on AlphaFold3
Results follow the same trend as AF2 Mode 3. The stronger the folding baseline, the smaller the performance gain. Improvements are mainly in pLDDT and RMSD, indicating our enhanced MSA primarily improves local structural regions while preserving the global protein architecture.
C.3 Folding Enhancement on Average Proteins
To probe the effectiveness of PLAME on average proteins, we firstly build a dataset from PDB validation set with 36 proteins. These protein MSAs don’t have sequence similarity over 90% compared to the PLAME training set. We randomly employ 32 MSAs for each protein and augment them with designed MSAs after HiFiAD filtering. The results are shown in Table 4.
pLDDT GDT TMscore RMSD LDDT pTM
AF2 MSA 83.156 0.767 0.785 5.243 0.753 0.718
PLAME 83.328 0.775 0.795 5.028 0.757 0.723
Table 4: Comparison of folding enhancement on average proteins
From the experimental results, the effects of augmentation align with our initial assumptions, demonstrating modest improvements. While the overall topological structure remains unchanged, minor adjustments can be observed in the structural details. As reported in MSAGPT, performance gains approach saturation between 16 and 32 augmentations. The relatively small improvements observed when applying our method to the average protein MSA can be attributed to the fact that these original MSAs already provide sufficient evolutionary information to AlphaFold2’s MSA Transformer, thus limiting the potential impact of additional augmentation.
C.4 Further Ablation on MSA Filtering
We further validate the effectiveness of filtered high-quality MSAs by comparing the performance with the more randomly selected MSAs (64 for each protein).
pLDDT GDT TMscore RMSD LDDT pTM
More Random MSAs 63.620 0.512 0.533 12.692 0.563 0.473
HiFiAD 66.349 0.534 0.553 11.755 0.581 0.506
Table 5: Comparison of folding enhancement based on different filterings.
From Table 5 and 2, We can observe a slight performance enhancement compared to Random-16 filtering approach according to pLDDT and LDDT. Conversely, the performance on global metric decreases. From the results, more co-evolutionary information may lead to better local geometric conformation, but it will disturb the modeling of the global conformations due to the bias during generation.
C.5 Failure Case Analysis
Other than analyzing successful cases, we analyzed four representative failure cases (3bog_B, 7sxb_A, 8gzu_AN, 8gzu_T3) with the largest performance drops, which includes three zero-shot and one few-shot examples. From the detailed results, we observe a clear mismatch between global metric, including GDT, TMScore, and RMSD, and local metric, including pLDDT, LDDT, and pTM on 3bog_B and 8gzu_T3. It is consistent with the metric discrepancies we observed in the main experiment.
pLDDT GDT TMscore RMSD LDDT pTM
AF2 MSA
3bog_B 41.493 0.150 0.130 22.443 0.148 0.129
7sxb_A 84.931 0.739 0.757 2.559 0.661 0.753
8gzu_AN 58.189 0.390 0.488 17.630 0.700 0.406
8gzu_T3 59.533 0.591 0.668 14.030 0.659 0.597
PLAME
3bog_B 32.918 0.169 0.148 17.522 0.158 0.118
7sxb_A 53.956 0.358 0.358 9.988 0.369 0.359
8gzu_AN 51.542 0.393 0.491 17.238 0.513 0.414
8gzu_T3 55.169 0.377 0.480 20.930 0.691 0.394
Table 6: Comparison of folding enhancement on failure cases.
Among the visualized MSA cases, we observed that generated MSAs contained extremely similar sequences (>90% similarity). Specifically, these high-similarity sequences caused all sites to appear more conserved, resulting in a lack of covariation patterns necessary for AlphaFold2 to infer structural contacts. This pattern was evident across all four cases. Notably, for 3bog_B and 8gzu_T3, the generated high-similarity MSAs further enhanced the conservation of already conserved regions, which consequently led to improvements in global metrics.
C.6 De novo Protein Folding Enhancement
We conduct further experiments on De Novo protein cases, where almost of them are orphan. Examples of de novo proteins include 8SK7 (RFDiffusion), 8TNM/8TNO (Chroma), and 8CYK (ProteinMPNN). We followed the same augmentation pattern as the main experiment.
pLDDT GDT TMscore RMSD LDDT pTM
AF2 MSA 89.27 0.886 0.904 1.658 0.781 0.800
HiFiAD 88.33 0.924 0.940 1.483 0.824 0.800
Table 7: Comparison of folding enhancement on de novo proteins.
From Table 7, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics. The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $>70\%$ similarity to the query sequences. This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset. Nevertheless, the diverse sampling strategy still effectively enhances the profile information of orphan proteins, resulting in substantial performance improvements. Furthermore, we visualized specific local regions where PLAME achieves superior alignment performance as measured by TMscore. Analysis revealed that across all augmented profiles, these high-performing local regions exhibit remarkable conservation, suggesting a strong correlation between sequence conservation patterns and structural alignment quality.
Figure 4: Comparison of structure enhancement of De Novo proteins.
Figure 5: Comparison of structure enhancement of De Novo proteins.
Appendix D Discussion
D.1 Limitations
Recent advancements in MSA generation models have shown promising results in enhancing protein folding predictions. However, several challenges remain to be addressed for broader applications and improved performance. 1) Limited quality by current model architectures, data constraints, and generation strategies, such as relying on small MSA prompts, hinders the overall richness and informativeness of the generated MSAs. Future methods should focus on constructing more expressive evolutionary latent spaces to better capture the complexity of protein sequence relationships and improve the informativeness of generated MSAs. 2) Distribution gaps still exist between the diversity and quality of generated MSAs and their natural counterparts, limiting their utility in broader applications. While current methods show potential in folding tasks, future models should focus on zero-shot generation capabilities to produce MSAs with distributions closer to natural MSAs, enabling broader applications such as conserved residue identification, mutation effect prediction, and functional annotation. 3) Assessing MSA quality remains an unresolved issue, as current evaluations primarily rely on downstream folding performance to infer quality. Developing direct and robust quality assessment metrics will be crucial for systematically evaluating and improving MSA generation methods, enabling the selection of high-quality MSAs for specific applications and paving the way for next-generation models with enhanced accuracy, broader applicability, and greater biological relevance.
D.2 Social Impact
PLAME advances multiple sequence alignment (MSA) generation for proteins, offering significant potential benefits for protein structure prediction and engineering applications. This capability could accelerate therapeutic protein development, enhance drug design processes, and facilitate the exploration of novel functional protein sequences. However, we acknowledge important limitations and risks associated with this technology. The model may occasionally generate hallucinated sequences that could mislead downstream protein design efforts if not properly validated. Furthermore, like many powerful biotechnology tools, there exists potential for misuse in designing harmful biological entities.
Appendix E Structure Comparison Visualization
Figure 6: Structure comparison visualization of 8ehb_F.
Figure 7: Structure comparison visualization of 8okh_B.
Figure 8: Structure comparison visualization of 8b4k_C.
Figure 9: Structure comparison visualization of 8fjf_A.
Figure 10: Structure comparison visualization of 8eoz_B.
Figure 11: Structure comparison visualization of 8okw_B.
Appendix F Augmented MSA Visualization
To provide an intuitive understanding of the MSAs generated by PLAME, we selected several representative cases for visualization. These cases demonstrate consistent improvements in folding accuracy compared to the MSAs provided by AF2 and cover a range of sequence lengths, including short ($<$100), medium (100-300), and long ($>$300) sequences, as well as cases under few-shot and zero-shot settings. For each visualization, the generated MSAs are highlighted with a black box. Additionally, the upper portion of each figure presents conservation information alongside the corresponding gap information. The protein information is provided in the left-top corner at each figure.
Figure 12: Augmented MSA visualization of 8ehb_F.
Figure 13: Augmented MSA visualization of 8okh_B.
Figure 14: Augmented MSA visualization of 8okw_B.
Figure 15: Augmented MSA visualization of 8fih_C.
Figure 16: Augmented MSA visualization of 7opb_D and T1119_D.
Figure 17: Augmented MSA visualization of 7xr1_A.
Figure 18: Augmented MSA visualization of 8e0n_F.
Appendix G Failure Case MSA Visualization
Figure 19: Failure Case MSA visualization of 7sxb_A.
Figure 20: Failure Case MSA visualization of 8gzu_T3.
Figure 21: Failure Case MSA visualization of 3bog_B.
Figure 22: Failure Case MSA visualization of 8gzu_AN.

### v2

@@
-PLAME: Leveraging Pretrained Language Models to Generate Enhanced Protein Multiple Sequence Alignments
+PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings
 Abstract
-Protein structure prediction is essential for drug discovery and understanding biological functions.
-While recent advancements like AlphaFold have achieved remarkable accuracy, most folding models rely heavily on multiple sequence alignments (MSAs) to boost prediction performance.
-This dependency limits their effectiveness on low-homology proteins and orphan proteins, where MSA information is sparse or unavailable.
-To address this limitation, we propose PLAME, a novel MSA design model that leverages evolutionary embeddings from pretrained protein language models.
-Unlike existing methods, PLAME introduces pretrained representations to enhance evolutionary information and employs a conservation-diversity loss to enhance generation quality.
-Additionally, we propose a novel MSA selection method to effectively screen high-quality MSAs and improve folding performance.
-We also propose a sequence quality assessment metric that provides an orthogonal perspective to evaluate MSA quality.
-On the AlphaFold2 benchmark of low-homology and orphan proteins, PLAME achieves state-of-the-art performance in folding enhancement and sequence quality assessment, with consistent improvements demonstrated on AlphaFold3.
-Ablation studies validate the effectiveness of the MSA selection method, while extensive case studies on various protein types provide insights into the relationship between AlphaFold’s prediction quality and MSA characteristics.
-Furthermore, we demonstrate that PLAME can serve as an adapter achieving AlphaFold2-level accuracy with the ESMFold’s inference speed.
+Protein structure prediction often hinges on multiple sequence alignments (MSAs), which underperform on low-homology and orphan proteins.
+We introduce PLAME, a lightweight MSA design framework that leverages evolutionary embeddings from pretrained protein language models to generate MSAs that better support downstream folding.
+PLAME couples these embeddings with a conservation–diversity loss that balances agreement on conserved positions with coverage of plausible sequence variation.
+Beyond generation, we develop (i) an MSA selection strategy to filter high-quality candidates and (ii) a sequence-quality metric that is complementary to depth-based measures and predictive of folding gains.
+On AlphaFold2 low-homology/orphan benchmarks, PLAME delivers state-of-the-art improvements in structure accuracy (e.g., lDDT/TM-score), with consistent gains when paired with AlphaFold3.
+Ablations isolate the benefits of the selection strategy, and case studies elucidate how MSA characteristics shape AlphaFold confidence and error modes.
+Finally, we show PLAME functions as a lightweight adapter, enabling ESMFold to approach AlphaFold2-level accuracy while retaining ESMFold-like inference speed.
+PLAME thus provides a practical path to high-quality folding for proteins lacking strong evolutionary neighbors.
 1 Introduction
-Understanding complex and highly variable protein structures is a cornerstone of drug design, enabling breakthroughs in target identification, validation, and drug-target interaction studies baker2001protein ; khoury2014protein ; jumper2021alphafold .
-Recent advancements, such as AlphaFold, have revolutionized structural biology by achieving near-experimental accuracy for most protein structures and their complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate .
-However, many of these state-of-the-art models rely heavily on evolutionary information encoded in Multiple Sequence Alignments (MSAs) as input jumper2021alphafold ; abramson2024accurate , making their predictive performance highly dependent on the quality of the MSAs.
-This dependency poses significant challenges for low-homology protein families and orphan proteins, where evolutionary information is sparse or unavailable kwon2021assessment ; webb2016comparative .
-Traditional approaches enhancing low-homology and orphan protein structure prediction include physics-based and template-based methods.
-Physics-based methods simulate energy functions but are computationally expensive and limited by approximations rohl2004rosetta ; cornell1995amber .
-Template-based methods rely on sequence homology detection hildebrand2009fast ; finn2011hmmer , making them unsuitable for orphan proteins where evolutionary information is scarce.
-These limitations have motivated the development of data-driven approaches to improve MSA generation.
-Figure 1: Taxonomy of existing MSA designers.
-Previous methods capture discrete-space distribution of MSAs, including sequence inpainting and prompt-based generation.
-PLAME serves as an evolutionary-space-inspired de novo generation method independent of sequence prompt.
-Recent MSA generation methods can be categorized into sequence space inpainting and prompt-based conditional generation approaches (Figure 1).
-Sequence inpainting methods like MSA Generator and EvoGen operate within sequence space by completing partial MSAs based on patterns from existing data zhang2023enhancing ; zhang2022few , aiming to reconstruct evolutionary constraints.
-Prompt-based conditional generation methods, such as MSAGPT and EvoDiff, utilize pretrained models conditioned on MSA prompts to generate new sequences chen2024msagpt ; alamdari2023protein .
-These AI-based approaches have shown potential in augmenting low-homology MSAs and enhancing structure prediction.
-As an alternative approach, models like ESMFold construct implicit evolutionary representations directly from single sequences using pretrained protein language models lin2023evolutionary .
-While this strategy circumvents the need for MSAs, these methods cannot utilize structural templates or enhanced homology information, leading to limited folding accuracy.
-Despite recent advancements, MSA generation methods targeting the enhancement of protein folding still face two major challenges.
-First, these models heavily rely on existing MSA databases, which are biased toward well-represented protein families and provide limited data for low-homology or orphan proteins.
-This reliance hinders their ability to construct evolutionary spaces that capture the unique constraints of such proteins, reducing their effectiveness in structure prediction.
-Second, there is a significant gap in understanding how MSA characteristics influence folding model performance.
-Current methods lack efficient evaluation metrics directly tied to folding performance, creating a disconnect between MSA generation objectives and structural accuracy.
-Additionally, the response of different protein types to MSA quality varies considerably, yet remains insufficiently studied.
-While approaches like MSAGPT have attempted to integrate structure prediction feedback through fine-tuning chen2024msagpt , their high computational cost limits scalability.
-A more comprehensive understanding of the MSA-folding relationship would enable more targeted optimization strategies.
-To address these limitations, we propose PLAME, a PLm Aligner for MSA Enhancement designed for low-homology protein structure enhancement. Our key contributions are as follows:
-  • Enriched evolutionary spaces: PLAME leverages pretrained protein language models to incorporate evolutionary embeddings as input, providing richer evolutionary information than sparse MSAs for low-homology and orphan proteins wang2024maape ; mcwhite2023leveraging ; hong2024fast .
-Furthermore, we introduce a novel conservation-diversity loss, which increases model sensitivity to conserved regions while maintaining sequence diversity, enabling the generation of biologically plausible MSAs optimized for high-quality alignments.
-  • Comprehensive MSA quality exploration: We propose a novel sequence quality assessment metric that provides a new perspective for evaluating MSA quality and its relationship to folding performance.
-Based on these insights, we develop an MSA selection method called High-Fidelity Appropriate Diversity (HiFiAD) that improves folding accuracy by effectively screening MSAs that balance sequence fidelity and diversity.
-This approach offers both analytical understanding and practical improvement for protein structure prediction.
-  • Extensive experimental validation: Our experimental results demonstrate the substantial enhancement on challenging low-homology and orphan protein datasets when evaluated with both AlphaFold2 and AlphaFold3 folding frameworks.
-Ablation studies confirm the effectiveness of the conservation-diversity loss and HiFiAD selection method in improving both MSA quality and folding performance.
-Additionally, we conduct comprehensive experiments on multiple protein systems, including general and de novo proteins, providing deeper insights into the relationship between MSA characteristics and folding performance.
+Understanding complex and variable protein structures is central to target identification, validation, and drug–target interaction studies in drug design baker2001protein ; khoury2014protein ; jumper2021alphafold .
+Recent advances such as AlphaFold have transformed structural biology, achieving near-experimental accuracy across a wide range of proteins and complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate .
+However, most state-of-the-art folding pipelines depend critically on evolutionary information encoded in Multiple Sequence Alignments (MSAs) jumper2021alphafold ; abramson2024accurate .
+As a consequence, their accuracy tracks the quality and depth of the available MSAs, leading to failure modes for low-homology families and orphan proteins where evolutionary neighbors are sparse or absent kwon2021assessment ; webb2016comparative .
+In practice, while MSA depth (e.g., effective sequence count) often correlates with predicted confidence (such as pLDDT), this relationship becomes highly unstable in sparse regimes where a few noisy or misaligned sequences can dominate the signal.
+Historically, two classes of techniques have been used to mitigate weak homology.
+Physics-based modeling seeks low-energy conformations under hand-crafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape rohl2004rosetta ; cornell1995amber .
+Template-based approaches leverage homology detection and profile–profile alignment to transfer structural priors from known folds hildebrand2009fast ; finn2011hmmer , yet degrade in regimes with poor evolutionary signal and are thus ill-suited to orphan proteins.
+These limitations have motivated data-driven strategies focused on improving the MSA itself rather than only the downstream folding network.
+Recent MSA design methods fall broadly into two lines (Figure 1).
+Sequence-space inpainting approaches (e.g., MSA Generator, EvoGen) complete or augment partial alignments by learning patterns directly in the discrete sequence space, aiming to reconstruct evolutionary constraints from observed MSAs zhang2023enhancing ; zhang2022few .
+Prompt-based conditional generation (e.g., MSAGPT, EvoDiff) uses pretrained models conditioned on MSA-style prompts to synthesize additional sequences chen2024msagpt ; alamdari2023protein .
+These approaches can deepen alignments and improve folding when homologs exist.
+An orthogonal line of work bypasses explicit MSAs by constructing implicit evolutionary representations from single sequences via large protein language models, as in ESMFold lin2023evolutionary .
+While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which can cap ultimate folding accuracy in challenging regimes.
+Figure 1: Taxonomy of MSA designers.
+Prior work models the discrete MSA distribution via sequence inpainting or prompt-based generation.
+PLAME operates in an evolutionary embedding space to de novo generate MSAs without requiring sequence prompts.
+Despite progress, two gaps remain for MSA design aimed at improving folding.
+(i) Biased supervision: Methods trained primarily on extant MSA databases inherit coverage biases toward well-studied families, offering limited guidance for low-homology and orphan proteins.
+This restricts the learned evolutionary manifold and hampers generalization exactly where better MSAs are most needed.
+(ii) Weak alignment–folding linkage: There is limited understanding—and few lightweight metrics—connecting MSA characteristics to downstream folding outcomes.
+As a result, generation objectives (e.g., likelihood in sequence space) can be misaligned with what most improves structure accuracy.
+While fine-tuning folding models with generated data (e.g., MSAGPT) partially closes the loop chen2024msagpt , the compute cost is high and does not yield general, model-agnostic criteria for selecting helpful MSAs.
+Closing this loop by retraining or adapting a folding model for each candidate MSA is prohibitive at scale, preventing iterative design-and-select cycles across large target sets.
+We introduce PLAME (PLm Aligner for MSA Enhancement), a lightweight MSA designer targeted at low-homology and orphan proteins.
+PLAME departs from discrete sequence modeling and instead generates in an evolutionary embedding space derived from pretrained protein language models, then decodes to MSAs for downstream folding.
+This design enables PLAME to synthesize plausible evolutionary neighborhoods even when observed homologs are scarce, while remaining compatible with template-using pipelines.
+  • Enriched evolutionary spaces.
+PLAME conditions on evolutionary embeddings from pretrained protein language models wang2024maape ; mcwhite2023leveraging ; hong2024fast , providing richer contextual signals than sparse MSAs for low-homology targets.
+We introduce a conservation–diversity loss that explicitly balances agreement at conserved positions with appropriate sequence variability, steering generation toward biologically plausible and alignment-friendly MSAs.
+  • Linking alignment to folding.
+We propose a complementary sequence-quality assessment metric that captures alignment properties predictive of folding gains, and we operationalize these insights with HiFiAD (High-Fidelity Appropriate Diversity), an MSA selection strategy that filters candidates to maximize downstream structure accuracy.
+  • Broad, model-agnostic validation.
+Across challenging low-homology and orphan datasets, PLAME improves folding accuracy with both AlphaFold2 and AlphaFold3.
+Ablations isolate the effects of the conservation–diversity loss and HiFiAD on MSA quality and folding outcomes, and case studies spanning general and de novo proteins elucidate how specific MSA characteristics modulate AlphaFold confidence and error modes.
 Figure 2: Overview of PLAME framework.
@@
 Here, $\mathbf{H}_{r}$ encodes the evolutionary space, facilitating cross-attention during decoding. The cross-row attention (Row-Attn) is defined as:
-[…]
+$\text{Row-Attn}(\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r})=\text{softmax}\left(\frac{\mathbf{Q}_{r}\mathbf{K}_{r}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{r}.$ (5)
 where $\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r}$ denote the Query, Key, and Value matrices, respectively.
@@
 The column attention is computed as:
-[…]
+$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}$ (7)
 Generation & Inference
@@
 The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position.
-These weights are normalized to the range $[1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation.
+These weights are normalized to the range $1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation.
 Specifically,
 $w_{l}=1+\delta\cdot\frac{\text{freq}_{l}-\text{min}(\text{freq})}{\text{max}(\text{freq})-\text{min}(\text{freq})}.$ (11)
+where freq denotes the residue-frequency of 20 types of amino acids.
 During model training, we apply $\delta=0.5$, assigning higher weights to conserved positions and lower weights to less conserved ones.
@@
 $S_{\text{BLOSUM}}(m_{i},s)=\sum_{j=1}^{L}B(s_{j},m_{ij}),\quad\forall m_{i}\in M,$ (16)
-where $B$ is the BLOSUM matrix (BLOSUM62 in our case).
+where $B$ is the BLOSUM matrix (BLOSUM62 in our case); $s_{j}$ represents the j-th amino acid of sequence S with length L, and $m_{ij}$ refers to the j-th position of the i-th MSA sequence with length L.
 For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution.
@@
 ESMFold 66.26 62.62 0.6 0.53 0.6 0.57 9.58 12.04 0.62 0.59 / /
-Mode1
+[…]
 AF2 MSA 60.07 62.14 0.50 0.52 0.50 0.57 12.34 12.16 0.54 0.58 0.44 0.49
@@
 PLAME 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
-Mode2
+[…]
 AF2 MSA 66.56 66.32 0.51 0.55 0.52 0.60 12.06 11.84 0.55 0.61 / /
@@
 PLAME 67.77 67.48 0.53 0.55 0.54 0.60 12.62 11.90 0.57 0.61 / /
-Mode3
+[…]
 AF2 MSA 70.31 69.61 0.57 0.60 0.57 0.64 10.53 10.24 0.60 0.65 / /
@@
 For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein .
-Any $>90\%$ redundancy between the test cases and training dataset is eliminated.
+Any $90\%$ redundancy between the test cases and training dataset is eliminated.
 4.1 PLAME as High-quality MSA Designer
@@
 We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics.
-Results are shown in Table 1.
+Results are shown in Table 3.4.
 We analyze the quality of MSAs generated by different methods and their performance in various scenarios.
@@
 Lemma 1.
-[…]
+Let $P(l,a)$ be the empirical amino–acid distribution for residue $a\!\in\!\mathcal{A}$, and let $Q_{\theta}(l,a)$ denote the model distribution at the residue (i.e.
+the conditional probability $p_{\theta}(a\mid y_{<l})$ after taking expectation over prefixes).
+Assign each column a weight $w_{l}\!\in[\,1-\delta,\,1+\delta\,]$ obtained from its conservation score.
+Then PCE loss directs optimization preferentially toward conserved positions by minimizing a weighted KL divergence and scaling gradient magnitudes in proportion to $w_{l}$.
 Proof.
@@
 $\mathbb{E}[\mathcal{L}_{\text{CE}}]=-\sum_{l=1}^{L}\sum_{a\in\mathcal{A}}P(l,a)\,\log Q_{\theta}(l,a).$ (17)
-Re-expressing each column term as $-\sum_{a}P\log Q=H\!\bigl{(}P(l,\!\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}$, we obtain
+Re-expressing each column term as $\sum_{a}P\log Q=H\!\bigl{(}P(l,\!\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}$, we obtain
 $\mathbb{E}[\mathcal{L}_{\text{CE}}]=\sum_{l=1}^{L}\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}+\sum_{l=1}^{L}H\!\bigl{(}P(l,\!\cdot)\bigr{)}.$ (18)
@@
 Based on the fact that $\theta^{\star}_{\mathrm{PCE}}$ minimizes $D_{w}$, denote $\Delta_{w}:=D_{w}(\theta^{\star}_{\mathrm{CE}})-D_{w}(\theta^{\star}_{\mathrm{PCE}})\geq 0$.
-By applying ()($\ast$)( ∗ ) to both optimal parameters and subtracting, we obtain:
+By applying $(\ast)$ to both optimal parameters and subtracting, we obtain:
 $(1-\delta)\Bigl{[}D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{CE}})-D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{PCE}})\Bigr{]}\;\leq\;\frac{\Delta_{w}}{L}.$ (30)
-Since $\Delta_{w}\geq 0$ and $01-\delta>$ 0; it is strictly positive whenever $\Delta_{w}>0$, Therefore,
+Since $\Delta_{w}\geq 0$ and $1-\delta>0$; it is strictly positive whenever $\Delta_{w}>0$, Therefore,
 $D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{PCE}})\;\leq\;D_{\text{KL}}^{\text{avg}}(\theta^{\star}_{\mathrm{CE}}),$ (31)
@@
 which is exactly the optima $Q_{\alpha}^{\star}(l,\cdot)$ mentioned earlier.
-Since $10<\tau<$ 1, this transformation always increases entropy unless $P$ is already uniform:
+Since $0<\tau<1$, this transformation always increases entropy unless $P$ is already uniform:
 $H\!\bigl{(}P(l,\cdot)\bigr{)}\;\leq\;H\!\bigl{(}Q_{\alpha}^{\star}(l,\cdot)\bigr{)}\;\leq\;\log|\mathcal{A}|.$ (37)
@@
 Sampling was executed on an A40 GPU.
-Appendix C Extensive Experimental Results
-C.1 Case Study on Successful Designs
-To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structrure visualization in Appendix F.
+Appendix C Extensive Benchmark Results
+C.1 Comparison to AI-based MSA Retrieval Methods
+We further compare the folding enhancement by MSAs retrieved by AI-based approaches [19].
+Without filtering, DHR’s performance is generally higher than AF2 MSA, EvoDiff, and MSAGPT, demonstrating the effectiveness of applying evolutionary embeddings like ESM2.
+Zero/Few pLDDT GDT TMscore RMSD LDDT
+DHR Zero 63.64 0.51 0.52 12.04 0.55
+DHR Few 62.60 0.52 0.57 11.92 0.59
+PLAME Zero 66.54 0.53 0.53 11.48 0.57
+PLAME Few 66.08 0.54 0.58 12.14 0.60
+Table 3: Comparison of folding enhancement to AI-based MSA retrieval method.
+C.2 Comparison on Inference Speed and Memory Usage
+To further demonstrate PLAME’s efficiency, we calculated the inference time and memory cost of each method.
+We used ENZYME 1.2.1.50 (EC Number) with length 488 as the test case.
+The results show that PLAME achieved the fastest speed among all AI-based methods while consuming only 4.5GB of memory.
+The processing speed is comparable to traditional methods like MMSeq2 and AI-based retrieval methods like DHR.
+Compared to retrieval-based methods, PLAME does not require downloading or building databases in advance, nor does it need preprocessing steps.
+This makes it more lightweight and efficient for deployment.
+Method Time per MSA (s) GPU Memory (Gb)
+PLAME 0.10 4.5
+DHR 0.16 + 358.61 (Alignment) 1.9
+MMSeq2 0.48 0.0
+MSAGPT 62.46 41.6
+EvoDiff 478.24 4.0
+Table 4: Comparison on inference speed and memory.
+C.3 Ablation on HiFiAD Filtering
+To demonstrate the effectiveness of HiFiAD, we conducted ablation experiments on filtering methods for other baselines.
+Among the benchmarks, PLAME maintains SOTA performance, and DHR-HiFiAD and PLAME show comparable performance across multiple metrics.
+Several metrics of EvoDiff, MSAGPT, and DHR (AI-based MSA retrieval methods) all demonstrate improvement after HiFiAD filtering compared to their original filtering methods, which directly proves the benefit of HiFiAD rules for MSA in AF-series structure prediction.
+This insight can help future work develop higher-quality MSA sequences.
+Zero/Few pLDDT GDT TMscore RMSD LDDT
+DHR-HiFiAD Zero 66.01 0.53 0.53 11.48 0.57
+DHR-HiFiAD Few 66.08 0.55 0.60 12.14 0.60
+EvoDiff-HifiAD Zero 58.24 0.46 0.46 13.74 0.51
+EvoDiff-HifiAD Few 60.89 0.49 0.54 12.39 0.56
+MSAGPT-HifiAD Zero 60.16 0.48 0.48 12.54 0.53
+MSAGPT-HifiAD Few 62.63 0.52 0.57 12.18 0.59
+Table 5: Ablation on HiFiAD filtering.
+C.4 Ablation on Protein Length
+We listed the performance of PLAME and AF2 MSA in different length ranges, and found that PLAME shows overall improvement, with the largest improvement in the 100-300 range, followed by the >>300 and <<100 ranges.
+We believe this is due to our MSA training data being mainly concentrated in the 100-300 length range.
+Length Range pLDDT GDT TMscore RMSD LDDT
+AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
+AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
+AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
+PLAME <<100 74.12 0.63 0.52 7.49 0.61
+PLAME 100-300 65.55 0.53 0.58 11.58 0.58
+PLAME >>300 58.31 0.45 0.53 16.16 0.54
+Table 6: Ablation on protein length.
+Appendix D Extensive Case Studies
+D.1 Case Study on Successful Designs
+To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix G.
 We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.
@@
 These represent an area requiring further improvement.
-C.2 AlphaFold3 Folding Experiment
+D.2 AlphaFold3 Folding Experiment
 To further validate PLAME is an universal MSA generation framework for different folding software, we conduct extensive experiment on AlphaFold3 [5].
@@
 Improvements are mainly in pLDDT and RMSD, indicating our enhanced MSA primarily improves local structural regions while preserving the global protein architecture.
-C.3 Folding Enhancement on Average Proteins
+D.3 Folding Enhancement on Average Proteins
 To probe the effectiveness of PLAME on average proteins, we firstly build a dataset from PDB validation set with 36 proteins.
@@
 The relatively small improvements observed when applying our method to the average protein MSA can be attributed to the fact that these original MSAs already provide sufficient evolutionary information to AlphaFold2’s MSA Transformer, thus limiting the potential impact of additional augmentation.
-C.4 Further Ablation on MSA Filtering
+D.4 Further Ablation on MSA Filtering
 We further validate the effectiveness of filtered high-quality MSAs by comparing the performance with the more randomly selected MSAs (64 for each protein).
@@
 From the results, more co-evolutionary information may lead to better local geometric conformation, but it will disturb the modeling of the global conformations due to the bias during generation.
-C.5 Failure Case Analysis
+D.5 Failure Case Analysis
 Other than analyzing successful cases, we analyzed four representative failure cases (3bog_B, 7sxb_A, 8gzu_AN, 8gzu_T3) with the largest performance drops, which includes three zero-shot and one few-shot examples.
@@
 pLDDT GDT TMscore RMSD LDDT pTM
-AF2 MSA
+[…]
 3bog_B 41.493 0.150 0.130 22.443 0.148 0.129
@@
 8gzu_T3 59.533 0.591 0.668 14.030 0.659 0.597
-PLAME
+[…]
 3bog_B 32.918 0.169 0.148 17.522 0.158 0.118
@@
 Notably, for 3bog_B and 8gzu_T3, the generated high-similarity MSAs further enhanced the conservation of already conserved regions, which consequently led to improvements in global metrics.
-C.6 De novo Protein Folding Enhancement
+D.6 De novo Protein Folding Enhancement
 We conduct further experiments on De Novo protein cases, where almost of them are orphan.
@@
 From Table 11, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics.
-The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $>70\%$ similarity to the query sequences.
+The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $70\%$ similarity to the query sequences.
 This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset.
@@
 Figure 5: Comparison of structure enhancement of De Novo proteins.
-Appendix D Discussion
-D.1 Limitations
+Appendix E Discussion
+E.1 Limitations
 Recent advancements in MSA generation models have shown promising results in enhancing protein folding predictions.
@@
 Developing direct and robust quality assessment metrics will be crucial for systematically evaluating and improving MSA generation methods, enabling the selection of high-quality MSAs for specific applications and paving the way for next-generation models with enhanced accuracy, broader applicability, and greater biological relevance.
-D.2 Social Impact
+E.2 Social Impact
 PLAME advances multiple sequence alignment (MSA) generation for proteins, offering significant potential benefits for protein structure prediction and engineering applications.
@@
 Furthermore, like many powerful biotechnology tools, there exists potential for misuse in designing harmful biological entities.
-Appendix E Structure Comparison Visualization
+Appendix F Structure Comparison Visualization
 Figure 6: Structure comparison visualization of 8ehb_F.
@@
 Figure 11: Structure comparison visualization of 8okw_B.
-Appendix F Augmented MSA Visualization
+Appendix G Augmented MSA Visualization
 To provide an intuitive understanding of the MSAs generated by PLAME, we selected several representative cases for visualization.
-These cases demonstrate consistent improvements in folding accuracy compared to the MSAs provided by AF2 and cover a range of sequence lengths, including short ($<$100), medium (100-300), and long ($>$300) sequences, as well as cases under few-shot and zero-shot settings.
+These cases demonstrate consistent improvements in folding accuracy compared to the MSAs provided by AF2 and cover a range of sequence lengths, including short (<<100), medium (100-300), and long (>>300) sequences, as well as cases under few-shot and zero-shot settings.
 For each visualization, the generated MSAs are highlighted with a black box.
@@
 Figure 18: Augmented MSA visualization of 8e0n_F.
-Appendix G Failure Case MSA Visualization
+Appendix H Failure Case MSA Visualization
 Figure 19: Failure Case MSA visualization of 7sxb_A.

### v3

@@
-PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings
+Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings
 Abstract
@@
 1 Introduction
-Understanding complex and variable protein structures is central to target identification, validation, and drug–target interaction studies in drug design baker2001protein ; khoury2014protein ; jumper2021alphafold .
-Recent advances such as AlphaFold have transformed structural biology, achieving near-experimental accuracy across a wide range of proteins and complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate .
-However, most state-of-the-art folding pipelines depend critically on evolutionary information encoded in Multiple Sequence Alignments (MSAs) jumper2021alphafold ; abramson2024accurate .
-As a consequence, their accuracy tracks the quality and depth of the available MSAs, leading to failure modes for low-homology families and orphan proteins where evolutionary neighbors are sparse or absent kwon2021assessment ; webb2016comparative .
-In practice, while MSA depth (e.g., effective sequence count) often correlates with predicted confidence (such as pLDDT), this relationship becomes highly unstable in sparse regimes where a few noisy or misaligned sequences can dominate the signal.
-Historically, two classes of techniques have been used to mitigate weak homology.
-Physics-based modeling seeks low-energy conformations under hand-crafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape rohl2004rosetta ; cornell1995amber .
-Template-based approaches leverage homology detection and profile–profile alignment to transfer structural priors from known folds hildebrand2009fast ; finn2011hmmer , yet degrade in regimes with poor evolutionary signal and are thus ill-suited to orphan proteins.
-These limitations have motivated data-driven strategies focused on improving the MSA itself rather than only the downstream folding network.
-Recent MSA design methods fall broadly into two lines (Figure 1).
-Sequence-space inpainting approaches (e.g., MSA Generator, EvoGen) complete or augment partial alignments by learning patterns directly in the discrete sequence space, aiming to reconstruct evolutionary constraints from observed MSAs zhang2023enhancing ; zhang2022few .
-Prompt-based conditional generation (e.g., MSAGPT, EvoDiff) uses pretrained models conditioned on MSA-style prompts to synthesize additional sequences chen2024msagpt ; alamdari2023protein .
-These approaches can deepen alignments and improve folding when homologs exist.
-An orthogonal line of work bypasses explicit MSAs by constructing implicit evolutionary representations from single sequences via large protein language models, as in ESMFold lin2023evolutionary .
-While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which can cap ultimate folding accuracy in challenging regimes.
 Figure 1: Taxonomy of MSA designers.
-Prior work models the discrete MSA distribution via sequence inpainting or prompt-based generation.
-PLAME operates in an evolutionary embedding space to de novo generate MSAs without requiring sequence prompts.
-Despite progress, two gaps remain for MSA design aimed at improving folding.
-(i) Biased supervision: Methods trained primarily on extant MSA databases inherit coverage biases toward well-studied families, offering limited guidance for low-homology and orphan proteins.
-This restricts the learned evolutionary manifold and hampers generalization exactly where better MSAs are most needed.
-(ii) Weak alignment–folding linkage: There is limited understanding—and few lightweight metrics—connecting MSA characteristics to downstream folding outcomes.
-As a result, generation objectives (e.g., likelihood in sequence space) can be misaligned with what most improves structure accuracy.
-While fine-tuning folding models with generated data (e.g., MSAGPT) partially closes the loop chen2024msagpt , the compute cost is high and does not yield general, model-agnostic criteria for selecting helpful MSAs.
-Closing this loop by retraining or adapting a folding model for each candidate MSA is prohibitive at scale, preventing iterative design-and-select cycles across large target sets.
-We introduce PLAME (PLm Aligner for MSA Enhancement), a lightweight MSA designer targeted at low-homology and orphan proteins.
-PLAME departs from discrete sequence modeling and instead generates in an evolutionary embedding space derived from pretrained protein language models, then decodes to MSAs for downstream folding.
-This design enables PLAME to synthesize plausible evolutionary neighborhoods even when observed homologs are scarce, while remaining compatible with template-using pipelines.
-  • Enriched evolutionary spaces.
-PLAME conditions on evolutionary embeddings from pretrained protein language models wang2024maape ; mcwhite2023leveraging ; hong2024fast , providing richer contextual signals than sparse MSAs for low-homology targets.
-We introduce a conservation–diversity loss that explicitly balances agreement at conserved positions with appropriate sequence variability, steering generation toward biologically plausible and alignment-friendly MSAs.
-  • Linking alignment to folding.
-We propose a complementary sequence-quality assessment metric that captures alignment properties predictive of folding gains, and we operationalize these insights with HiFiAD (High-Fidelity Appropriate Diversity), an MSA selection strategy that filters candidates to maximize downstream structure accuracy.
-  • Broad, model-agnostic validation.
-Across challenging low-homology and orphan datasets, PLAME improves folding accuracy with both AlphaFold2 and AlphaFold3.
-Ablations isolate the effects of the conservation–diversity loss and HiFiAD on MSA quality and folding outcomes, and case studies spanning general and de novo proteins elucidate how specific MSA characteristics modulate AlphaFold confidence and error modes.
-Figure 2: Overview of PLAME framework.
-PLAME captures ESM-2 evolutionary representations, generating MSAs for augmenting the original MSAs.
-The augmented MSAs serve as the homology template for folding softwares for folding enhancement.
-In each block of the T5-architecture, additional row-attention and col-attention are applied to capture co-evolutionary information.
+Understanding complex and dynamic protein structures is fundamental to target identification, validation, and drug-target interaction studies in drug design (baker2001protein, ; khoury2014protein, ).
+Recent advances such as AlphaFold have revolutionized structural biology, achieving near-experimental accuracy across a broad spectrum of proteins and complexes (jumper2021alphafold, ; ahdritz2024openfold, ; abramson2024accurate, ).
+However, most state-of-the-art folding pipelines heavily rely on evolutionary information encoded within multiple sequence alignments (MSAs) (lin2023evolutionary, ; abramson2024accurate, ).
+Consequently, their accuracy is highly correlated with the quality and depth of available MSAs.
+This dependency creates failure modes in low-homology families and orphan proteins (those lacking or having few evolutionary neighbors) (kwon2021assessment, ; webb2016comparative, ), where even small amounts of noisy or misaligned sequences can dominate the signal.
+Historically, two primary classes of techniques have been developed to address weak homology.
+Physics-based modeling searches for low-energy conformations in energy space through handcrafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape (rohl2004rosetta, ; cornell1995amber, ).
+Template-based methods leverage homology detection and profile-profile alignment to transfer structural priors from known folds to novel sequences (hildebrand2009fast, ; finn2011hmmer, ), but suffer degraded performance in the absence of evolutionary signals, making them unsuitable for orphan proteins.
+These limitations have motivated a shift toward data-driven strategies that focus on improving the MSA itself rather than solely the downstream folding networks.
+Recent MSA design approaches can be broadly categorized into two paradigms (Figure 1).
+Sequence-space inpainting methods (e.g., MSA Generator, EvoGen) directly learn patterns in discrete sequence space to augment partial alignments, aiming to reconstruct evolutionary constraints from existing MSAs (zhang2023enhancing, ; zhang2022few, ).
+Prompt-based conditional generation approaches (e.g., MSAGPT, EvoDiff) utilize pre-trained models to synthesize additional sequences under MSA-style prompts (chen2024msagpt, ; alamdari2023protein, ).
+These methods can deepen alignments and improve folding accuracy when homologous sequences exist.
+An orthogonal line of research bypasses explicit MSA construction by building implicit evolutionary representations from single sequences through large protein language models (PLMs), as demonstrated by ESMFold (lin2023evolutionary, ).
+While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which may limit ultimate folding accuracy in challenging scenarios.
+Despite existing progress, two critical gaps remain in structure prediction for low-homology proteins.
+(i) Supervision bias: Methods trained on existing MSA databases inherit biases toward well-studied families, limiting effectiveness for low-homology and orphan proteins.
+(ii) Weak alignment-folding correlation: Current approaches lack lightweight metrics linking MSA characteristics to folding outcomes.
+Sequence-based generative objectives may not align with factors that improve structural accuracy, while existing solutions like fine-tuning folding models (chen2024msagpt, ) are computationally expensive and lack universal applicability.
+In this study, we propose PLAME, motivated by the critical need to enhance structure prediction for low-homology proteins where traditional MSA-based approaches fail due to insufficient evolutionary signals.
+Our approach makes the following key contributions:
+  1. 1.
+    Embedding-space MSA generation with conservation-diversity optimization: Inspired by PLMs’ success in MSA-related tasks (hong2024fast, ; wang2024maape, ; mcwhite2023leveraging, ), we develop the first MSA designer that generates auto-regressively within the evolutionary embedding space of pre-trained PLMs rather than discrete sequences (Fig2).
+We further propose a novel conservation-diversity loss that captures conserved regions while extracting diverse variants from ESM embeddings with theoretical guarantee (AppendixA).
+The lightweight design enables PLAME to synthesize evolutionary neighborhoods even with scarce homologous sequences, achieving up to three orders of magnitude speedup while maintaining template compatibility (Table 4).
+  2. 2.
+    HiFiAD: A principled MSA quality assessment framework: To address the current weak alignment-folding correlation problem, we propose High-Fidelity Appropriate Diversity (HiFiAD), a lightweight algorithm for MSA filtering that simultaneously considers site-wise conservation and inter-MSA diversity.
+This provides the first model-agnostic, computationally efficient criterion for selecting high-quality alignments that directly correlate with improved folding outcomes.
+  3. 3.
+    Comprehensive validation across challenging scenarios: On challenging low-homology and orphan datasets, PLAME consistently improves folding accuracy in both AlphaFold2 and AlphaFold3, performing similarly to DHR (hong2024fast, ), AI-based MSA searching approach.
+In ablation studies, HiFiAD demonstrates performance gains across all baselines (Table4).
+Moreover, case studies on general and de novo proteins further demonstrate PLAME’s generalizability while providing novel perspectives on structure enhancement from an MSA design standpoint (Table8).
+PLAME offers new insights and possibilities for folding enhancement through principled MSA optimization.
 2 Related Works
@@
 Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches.
-Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding cornell1995amber ; brooks2009charmm .
-While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results karplus2002molecular ; freddolino2010challenges ; pande2010everything .
-Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures rohl2004rosetta ; hildebrand2009fast .
-These methods perform well with suitable templates but struggle with orphan proteins and low-homology families webb2016comparative ; baker2001protein .
-Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability jumper2021alphafold ; abramson2024accurate ; wu2022high .
+Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding (cornell1995amber, ; brooks2009charmm, ).
+While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results (karplus2002molecular, ; freddolino2010challenges, ; pande2010everything, ).
+Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures (rohl2004rosetta, ; hildebrand2009fast, ).
+These methods perform well with suitable templates but struggle with orphan proteins and low-homology families (webb2016comparative, ; baker2001protein, ).
+Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability (jumper2021alphafold, ; abramson2024accurate, ; wu2022high, ).
 Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
@@
 The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy.
-However, these methods often slow down inference despite their modest performance gains johnson2010hidden ; steinegger2017mmseqs2 ; zheng2024improving ; lee2024petabase .
+However, these methods often slow down inference despite their modest performance gains (johnson2010hidden, ; steinegger2017mmseqs2, ; zheng2024improving, ; lee2024petabase, ).
 The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency.
-However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy lin2023evolutionary ; mirdita2022colabfold .
+However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy (lin2023evolutionary, ; mirdita2022colabfold, ).
 The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families.
-While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability alamdari2023protein ; zhang2022few ; zhang2023enhancing ; chen2024msagpt .
+While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability (alamdari2023protein, ; zhang2022few, ; zhang2023enhancing, ; chen2024msagpt, ).
+Figure 2: Overview of PLAME framework.
+PLAME captures ESM-2 evolutionary representations, generating MSAs for augmenting the original MSAs.
+The augmented MSAs serve as the homology template for folding softwares for folding enhancement.
+In each block of the T5-architecture, additional row-attention and col-attention are applied to capture co-evolutionary information.
 3 Method
@@
 where the augmented MSAs are composed of original MSAs $\mathbf{M}$ and generated MSAs $\mathbf{M}^{\prime}$, denoted as $\mathbf{M}_{\text{aug}}=\{\mathbf{M},\mathbf{M}^{\prime}\}$.
-The quality of the enhanced structures is evaluated using several metrics, including RMSD, TM-score, and pLDDT.
+The quality of the enhanced structures is evaluated using several metrics, including RMSD, TM-score, and pLDDT (See details in Section4).
 The key to high-fidelity MSA generation lies in constructing an informative evolutionary distribution $\mathbf{z}_{\text{evo}}$, which serves as the foundation for generating augmented MSAs $\mathbf{M}_{\text{aug}}$.
@@
 3.2 Model Architecture
-PLAME employs an encoder-decoder transformer architecture similar to MSA Transformer rao2021msa , with adjustments to the T5 block structure vaswani2017attention .
-The encoder and decoder incorporate additional row-wise and column-wise attention mechanisms, designed to better capture evolutionary patterns in MSA data (detailed in Fig 2).
+PLAME employs an encoder-decoder transformer architecture similar to MSA Transformer (rao2021msa, ), with adjustments to the T5 block structure (vaswani2017attention, ).
+The encoder and decoder incorporate additional row-wise and column-wise attention mechanisms to better capture evolutionary patterns in MSA data (detailed in Fig 2), which is similarly applied in MSAGenerator (zhang2023enhancing, ) and MSAGPT (chen2024msagpt, ).
 Additional mechanisms are introduced as follows.
 Row Attention
-The row attention mechanism models inter-sequence dependencies within the evolutionary space, summarizing sequence relationships across the depth of the input representation.
-Given an input tensor $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{L\times N\times D\times h}$, where $L$ is the sequence length, $N$ is the number of MSAs, $D$ is the depth, and $h$ is the hidden dimension, a global representation $\mathbf{H}r$ is computed by averaging the hidden states along the depth dimension:
-$\mathbf{H}_{r}=\frac{1}{D}\sum_{d=1}^{D}\mathbf{H}_{\text{enc}}^{d},\quad\mathbf{H}_{r}\in\mathbb{R}^{L\times N\times h}.$ (4)
-Here, $\mathbf{H}_{r}$ encodes the evolutionary space, facilitating cross-attention during decoding. The cross-row attention (Row-Attn) is defined as:
+Row attention models inter-sequence dependencies by summarizing evolutionary relationships across MSA depth.
+Given input $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{L\times N\times D\times h}$, we compute a global representation by depth-wise averaging:
+$\mathbf{H}_{r}=\frac{1}{D}\sum_{d=1}^{D}\mathbf{H}_{\text{enc}}^{d}\in\mathbb{R}^{L\times N\times h},$ (4)
+where $\mathbf{H}_{r}$ encodes the evolutionary space for cross-attention during decoding:
 $\text{Row-Attn}(\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r})=\text{softmax}\left(\frac{\mathbf{Q}_{r}\mathbf{K}_{r}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{r}.$ (5)
-where $\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r}$ denote the Query, Key, and Value matrices, respectively.
 Column Attention
-Column attention captures conservation patterns across columns in the MSA, focusing on evolutionary signals at specific positions.
-To implement this mechanism, the input representation matrix $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{D\times N\times L\times h}$ is first transposed to $\mathbf{X}_{\text{dec}}^{\top}\in\mathbb{R}^{L\times N\times D\times h}$.
-The self and cross-column attention mechanisms operate on this transformed representation.
-For cross-column attention, the query, key, and value matrices $\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}}$ and their corresponding projection weights $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are defined as:
-$\mathbf{Q}_{\text{c}}=\mathbf{X}_{\text{dec}}^{\top}\mathbf{W}_{q},\quad\mathbf{K}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{k},\quad\mathbf{V}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{v}.$ (6)
-The column attention is computed as:
-$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}$ (7)
+Column attention captures positional conservation patterns across MSA columns.
+We transpose the decoder input $\mathbf{X}_{\text{dec}}\in\mathbb{R}^{D\times N\times L\times h}$ to $\mathbf{X}_{\text{dec}}^{\top}$ and compute cross-column attention with:
+$\mathbf{Q}_{\text{c}}=\mathbf{X}_{\text{dec}}^{\top}\mathbf{W}_{q},\quad\mathbf{K}_{\text{c}}$ $=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{k},\quad\mathbf{V}_{\text{c}}=\mathbf{H}_{\text{enc}}^{\top}\mathbf{W}_{v},$ (6)
+$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})$ $=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}.$ (7)
 Generation & Inference
-In our implementation, ESM-2 ($\mathbf{g}_{\phi}$) encodes the query sequence $\mathbf{s}$ into high-dimensional evolutionary embeddings, denoted as $\mathbf{H}_{\text{input}}$.
-These embeddings are processed by the encoder through $N$ layers of modified T5 blocks, iteratively refining contextualized representations:
+ESM2 (lin2023evolutionary, ) encodes the query sequence $\mathbf{s}$ into evolutionary embeddings $\mathbf{H}_{\text{input}}$. The encoder processes these through $N$ modified T5 layers:
 $\mathbf{H}_{\text{Enc}}^{(l)}=\mathbf{Enc}^{(l)}(\mathbf{H}^{(l-1)}),\quad l=1,\dots,N,\quad\mathbf{H}^{(0)}=\mathbf{H}_{r}.$ (8)
-The final encoder output, $\mathbf{H}^{(N)}$, captures comprehensive sequence context and is passed to the decoder for autoregressive generation.
-The decoder generates tokens sequentially, conditioned on the encoder output $\mathbf{H}_{\text{Enc}}^{(N)}$ and previously generated tokens:
+The decoder autoregressively generates tokens conditioned on encoder output and previous tokens:
 $\mathbf{y}_{t}=\mathbf{Dec}(\mathbf{y}_{<t},\mathbf{H}_{\text{Enc}}^{(N)}).$ (9)
-The T5 blocks with row and column attention enable the decoder to compute output embeddings, which are then passed through a softmax layer to produce token probabilities for the next position.
+Output embeddings are passed through softmax to produce token probabilities.
 3.3 Conservation-Diversity Training Loss
-We introduce a position-aware causal inference approach for diverse MSA generation, integrating a PSSM-Weighted Cross-Entropy Loss and a Diversity Regularization Loss to balance focus on conserved regions with sampling diversity.
-PSSM-Weighted Cross Entropy (PCE) Loss
-The PSSM-Weighted Cross Entropy Loss emphasizes accurate predictions in conserved regions of the MSA, which are critical for maintaining protein structure and function.
-For a single sequence, it is defined as:
+We propose a position-aware causal inference approach for diverse MSA generation, integrating a PSSM-Weighted Cross-Entropy (PCE) Loss and a DIversity REgularization (DIRE) Loss to balance focus on conserved regions with sampling diversity.
+PCE Loss
+The PCE Loss emphasizes accurate predictions in conserved regions of the MSA, which are critical for maintaining protein structure and function. For a single sequence, it is defined as:
 $\mathcal{L}_{\text{seq}}=-\sum_{l=1}^{L}w_{l}\cdot\log p(y_{l}\mid y_{<l}),$ (10)
-where $L$ is the sequence length, $y_{l}$ is the true label at position $l$, and $p(y_{l}\mid y_{<l})$ is the predicted probability of $y_{l}$.
-The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position.
+where $L$ denotes sequence length, $y_{l}$ denotes the amino acid at site $l$, and $p(y_{l}\mid y_{<l})$ denotes the predicted discrete probability distribution of $y_{l}$.
+The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) (henikoff1994position, ) and reflect the conservation level at each position.
 These weights are normalized to the range $1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation.
@@
 This loss emphasizes conserved regions while allowing flexibility in less conserved areas.
-Diversity Regularization (DIRE) Loss
-The Diversity Regularization Loss promotes biological diversity in MSAs by maximizing the entropy of predicted amino acid distributions.
-For a single sequence, the entropy at position $l$ is calculated as:
-$H_{l}=-\sum_{a\in\mathcal{A}}p(a\mid y_{<l})\log p(a\mid y_{<l}),$ (13)
-where $p(a\mid y_{<l})$ is the predicted probability of amino acid $a$ at position $l$, and $\mathcal{A}$ is the set of all amino acids.
-To encourage diversity across a batch of $N$ sequences, we compute the average entropy over all positions and sequences:
-$\mathcal{L}_{\text{diversity}}=-\frac{1}{N}\sum_{j=1}^{N}\frac{1}{L_{j}}\sum_{l=1}^{L_{j}}H_{l}^{(j)},$ (14)
-where $L_{j}$ is the length of sequence $j$, and $H_{l}^{(j)}$ is the entropy at position $l$ in sequence $j$.
-This loss encourages the model to capture the natural amino acid diversity in homologous sequences.
+DICE Loss
+The DIRE loss promotes sequence diversity by maximizing amino acid entropy:
+$\mathcal{L}_{\text{DIRE}}=-\frac{1}{N}\sum_{j=1}^{N}\frac{1}{L_{j}}\sum_{l=1}^{L_{j}}H_{l}^{(j)},$ (13)
+where $H_{l}^{(j)}=-\sum_{a\in\mathcal{A}}p(a\mid y_{<l})\log p(a\mid y_{<l})$ is the entropy at position $l$ in sequence $j$, and $\mathcal{A}$ is the set of all amino acids.
 Combined Loss Function
@@
 $\mathcal{L}=\alpha\cdot\mathcal{L}_{\text{PCE}}+(1-\alpha)\cdot\mathcal{L}_{\text{DIRE}},$ (14)
-where $\alpha\in[0,1]$ controls the trade-off between the two terms.
-We set $\alpha=0.9$ in our experiments, prioritizing conservation while maintaining sufficient diversity.
-This design ensures the generated MSAs accurately capture conserved regions and retain natural sequence variability.
-Our theoretical analysis in Appendix A demonstrates that PCE Loss enhances the model’s understanding of evolutionary information (MSA profile), while DIRE Loss functions as a regularizer to prevent neglect of variable regions.
+with $\alpha=0.9$ prioritizing conservation while maintaining variability.
+Our theoretical analysis in AppendixA demonstrates that PCE Loss enhances the model’s understanding of MSA profile, while DIRE Loss functions as a regularizer to prevent neglect of variable regions.
 3.4 MSA Selection Method – HiFiAD
-Existing MSA generation models struggle in ensuring the quality of generated MSAs.
-Different methods capture MSA distributions with varying effectiveness, and excessive concatenation of virtual MSAs can degrade the quality of the original MSA.
-This, in turn, restricts the ability of protein structure prediction software to interpret evolutionary distributions.
-While MSAGPT has explored MSA selection in its ablation studies, existing approaches lack a clear definition of MSA quality and fail to propose systematic selection methods based on sequence-level quality metrics chen2024msagpt .
-During MSA augmentation, virtual MSAs inevitably introduce noise, limiting the performance of folding software.
-A common issue with current models is the generation of over-conserved sequences that closely resemble the query sequence.
-When these sequences are over-concatenated, they distort the evolutionary distribution of the original MSA.
-This issue was highlighted in MSAGPT’s studies, which showed that performance plateaued once the number of generated MSAs exceeded 32.
-To address this, balancing recovery rate distributions by selecting sequences with both high and low recovery rates has proven more effective than focusing exclusively on one or the other.
-To address these issues, we propose HiFiAD to balance fidelity and diversity for MSA selection while maintaining efficiency.
-For orphan proteins without original MSAs (termed zero-shot cases), we select the top-$k$ sequences with the highest $S_{\text{BLOSUM}}$, along with sequences from the top and bottom $k/2$ of the recovery rate distribution.
-The BLOSUM substitution score is calculated as:
+Generated MSAs often contain noise that degrades folding performance. We propose HiFiAD to balance fidelity and diversity in MSA selection through principled quality metrics.
+HiFiAD addresses two key challenges: (i) over-conserved sequences that distort evolutionary distributions when over-concatenated, and (ii) lack of systematic quality assessment for generated MSAs, by combining sequence similarity (fidelity) with diversity to maintain balanced evolutionary signals.
+Given a query sequence $s$ and generated MSAs $M=\{m_{1},m_{2},\ldots,m_{n}\}$, we define:
 $S_{\text{BLOSUM}}(m_{i},s)$ $=\sum_{j=1}^{L}B(s_{j},m_{ij}),\quad\forall m_{i}\in M,$ (15)
-where $B$ is the BLOSUM matrix (BLOSUM62 in our case); $s_{j}$ represents the j-th amino acid of sequence S with length L, and $m_{ij}$ refers to the j-th position of the i-th MSA sequence with length L.
-For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution.
-Specifically, we cap the number of MSA sequences $N_{\text{aug}_{\text{max}}}$ at $\max(16,2N)$.
-If fewer than $N_{\text{aug}_{\text{max}}}$ virtual MSAs are available, we retain all generated MSAs to ensure sufficient diversity for downstream analysis.
+$R(m_{i},s)$ $=\frac{1}{L}\sum_{j=1}^{L}\mathbb{I}[s_{j}=m_{ij}],\quad\forall m_{i}\in M,$ (16)
+where $B$ is the BLOSUM62 matrix, $R(m_{i},s)$ is the recovery rate, and $\mathbb{I}[\cdot]$ is the indicator function.
+Zero-shot selection (Orphan proteins): Select top-$k$ sequences by $S_{\text{BLOSUM}}$ and sequences from top/bottom $k/2$ of recovery rate distribution, similar to the Static Diversity Strategy of MSAGPT.
+Few-shot selection (Low homology proteins): Limit augmented MSAs to $N_{\text{max}}=\max(16,2N_{\text{orig}})$ where $N_{\text{orig}}$ is the original MSA count.
+This design prevents evolutionary information distortion caused by excessive generated MSAs.
+4 Experiment
+Baselines
+To evaluate PLAME’s capability in generating high-fidelity and diverse MSAs, we compared it with several state-of-the-art AI-based MSA generation methods and AlphaFold2’s MSA pipeline (jumper2021alphafold, ).
+The baselines include AF2 MSA (johnson2010hidden, ), and open-source methods including EvoDiff and MSAGPT (chen2024msagpt, ; alamdari2023protein, ).
+Additionally, we include an MSA-free method, ESMFold (lin2023evolutionary, ), to evaluate the complementary benefits of explicit MSA enhancement versus implicit evolutionary modeling.
+Datasets
+For the training dataset, we used the PDB and UniClust30 subsets from the OpenProteinSet as our data source (ahdritz2024openproteinset, ).
+The pre-searched MSAs from OpenFold training were also included.
+We retained data with at least 64 MSA sequences.
+To avoid overlap with the test cases, we removed sequences with over 90% similarity by MMSeqs based on UniClust30 clustering results (mirdita2017uniclust, ; steinegger2017mmseqs2, ).
+This process yielded an initial dataset of 293,979 samples, which were split into training and validation sets with a 90:10 ratio.
+For the test dataset, we adopted the curated test cases from MSAGPT (chen2024msagpt, ), which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO (haas2018continuous, ), and PDB (berman2000protein, ).
+Any $90\%$ redundancy between the test cases and training dataset was eliminated.
+Evaluation
+Structural Assessment Metric We evaluate structure qualitywith local and global metrics.
+Local metrics include pLDDT (per-residue confidence) and LDDT (local distance difference test).
+Global metrics comprise GDT (global distance test), TM-Score (template modeling score) (zhang2005tm, ), pTM (predicted TM-score), and RMSD (root mean square deviation).
+AlphaFold2 Folding Modes To comprehensively assess MSA augmentation effectiveness, we evaluate three AF2 configurations with increasing computational complexity:
+Sequence Assessment Metric We employ four sequence-based metrics to quantify alignment fidelity and diversity:
+1) Conservation Score measures residue conservation at each position: $C_{i}=\text{Freq}_{\text{max}}(i)/N$, where $\text{Freq}_{\text{max}}(i)$ is the most frequent residue at position $i$ and $N$ is the sequence count.
+Higher scores indicate stronger evolutionary constraints.
+2) Gap Proportion quantifies alignment completeness: $G_{i}=G(i)/N$, where $G(i)$ counts gaps at position $i$. Lower values indicate better alignment quality.
+3) Substitution Compatibility evaluates evolutionary plausibility using BLOSUM62 scores $S_{\text{BLOSUM}}$ (Eq. 15). Higher scores reflect greater biological relevance.
+4) Alignment Entropy captures positional diversity via Shannon entropy:
+$H_{i}=-\sum_{r\in\{R_{i}\}}p(r)\log_{2}p(r)$ (17)
+where $\{R_{i}\}$ represents unique residues at position $i$ and $p(r)=\text{count}(r)/N$. Higher entropy indicates greater diversity; lower entropy suggests functional conservation.
 Table 1: Performance metrics across different modes and models.
@@
 Zero and Few indicate zero-shot (proteins without MSAs) and few-shot cases (proteins with existing MSAs), respectively.
-pLDDT GDT TMscore RMSD LDDT pTM
+pLDDT  ($\uparrow$) GDT  ($\uparrow$) TMscore  ($\uparrow$) RMSD($\downarrow$) LDDT  ($\uparrow$) pTM  ($\uparrow$)
 Zero Few Zero Few Zero Few Zero Few Zero Few Zero Few
@@
 MSAGPT 59.81 61.18 0.48 0.51 0.48 0.56 12.62 12.35 0.53 0.57 0.43 0.48
-PLAME 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
+DHR 63.64 62.60 0.51 0.52 0.52 0.57 12.04 11.92 0.55 0.59 / /
+[…]
 […]
@@
 MSAGPT 64.88 65.96 0.51 0.56 0.51 0.60 12.60 11.90 0.55 0.61 / /
-PLAME 67.77 67.48 0.53 0.55 0.54 0.60 12.62 11.90 0.57 0.61 / /
+[…]
 […]
@@
 MSAGPT 68.39 69.30 0.57 0.60 0.56 0.64 11.05 10.40 0.59 0.64 / /
-PLAME 71.50 70.48 0.58 0.59 0.58 0.64 11.41 10.62 0.60 0.64 / /
-4 Experiment
-Baselines
-To evaluate PLAME’s capability in generating high-fidelity and diverse MSAs, we compared it with several state-of-the-art AI-based MSA generation methods and AlphaFold2’s MSA pipeline jumper2021alphafold .
-The baselines include AF2 MSA, and open-source methods including EvoDiff and MSAGPT chen2024msagpt ; alamdari2023protein .
-Additionally, to assess the potential advantages of leveraging AlphaFold2’s knowledge base in protein structure prediction, we included ESMFold lin2023evolutionary , an MSA-free folding software, as reference.
-Evaluation Metric
-We apply the following metrics for structure assessment: 1) local metrics including pLDDT and LDDT; 2) global metrics including GDT, TM-Score, pTM, and RMSD.
-To comprehensively evaluate the effectiveness of augmented MSAs, we tested three AF2 folding mode, progressing from computationally efficient, basic settings to more comprehensive configurations: 1) Mode1: Folding using pTM-3 model without using templates; 2) Mode2: Folding using the default 5 models without using templates; 3) Mode3: Folding using the default 5 models with templates.
-To evaluate the quality of generated multiple sequence alignments (MSAs), we employ four sequence-based metrics to quantify alignment fidelity and diversity: 1) Conservation Score measures the degree of residue conservation at each alignment position.
-It is defined as $C_{i}=\frac{\text{Freq}_{\text{max}}(i)}{N}$, where $\text{Freq}_{\text{max}}(i)$ is the frequency of the most common residue at position $i$, and $N$ is the total number of sequences.
-A high score indicates strong conservation, often reflecting functional or evolutionary constraints, while a low score suggests positional variability.
-2) Gap Proportion quantifies the fraction of gaps at each alignment position, calculated as $G_{i}=\frac{G(i)}{N}$, where $G(i)$ is the number of gaps at position $i$.
-Lower values indicate better alignment quality with fewer missing data or alignment errors.
-3) Substitution Compatibility assesses the evolutionary plausibility of aligned residues using BLOSUM62 substitution scores $S_{\text{BLOSUM}}$ (Eq.
-16).
-Higher scores indicate evolutionary compatibility and alignment biological relevance.
-4) Alignment Entropy captures residue diversity at each position using Shannon entropy:
-$H_{i}=-\sum_{r\in\{R_{i}\}}p(r)\log_{2}p(r),$
-where $\{R_{i}\}$ is the set of unique residues at position $i$, and $p(r)=\frac{\text{count}(r)}{N}$ is the relative frequency of residue $r$.
-Higher entropy signifies greater positional diversity, indicating weaker functional constraints, whereas lower entropy denotes stronger conservation, often associated with functional or evolutionary significance.
-Datasets
-For the training dataset, we use the PDB and UniClust30 subsets from the open-source OpenProteinSet as our data source ahdritz2024openproteinset .
-The pre-searched MSAs from OpenFold training are also included.
-We only retain data with at least 64 MSA sequences.
-To avoid overlap with the test cases, we removed sequences with over 90% similarity using MMSeqs based on UniClust30 clustering results mirdita2017uniclust ; steinegger2017mmseqs2 .
-This process yields an initial dataset of 293,979 samples, which are split into training and validation sets with a 90:10 ratio.
-For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein .
-Any $90\%$ redundancy between the test cases and training dataset is eliminated.
-4.1 PLAME as High-quality MSA Designer
-General folding performance
-We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics.
-Results are shown in Table 3.4.
-We analyze the quality of MSAs generated by different methods and their performance in various scenarios.
-1) Performance Superiority: PLAME achieved state-of-the-art performance on most metrics across all modes.
-This demonstrates that PLAME-generated MSAs effectively enhance AF2 folding, particularly in scenarios involving low-homology or orphan proteins, where significant improvements in prediction accuracy were observed.
-In contrast, EvoDiff and MSAGPT struggled to produce high-quality MSAs under these conditions.
-By leveraging the evolutionary latent space provided by ESM-2, PLAME generated biologically meaningful virtual MSAs, outperforming baselines on key metrics such as pLDDT and RMSD.
-2) Diminishing Returns: As more advanced folding configurations were employed (e.g., with more powerful AF2 models or the inclusion of templates), the benefits of MSA augmentation gradually decreased.
-In Mode3, where templates were used, certain metrics exhibit performance decrease.
-This is likely because template information already captures much of the necessary sequence and evolutionary information, reducing the marginal utility of additional MSA augmentation.
-Furthermore, the noise introduced by low-quality MSAs can be amplified in stronger baselines, further impacting performance.
-3) Metric Discrepancies: The enhancement effects across different metrics were not entirely consistent.
-For instance, the trends for pLDDT and RMSD diverged.
-While pLDDT, as a confidence metric, reflects the model’s confidence in its predictions, RMSD measures the global deviation of the predicted structure from the ground truth.
-MSA augmentation improved local prediction quality (leading to higher pLDDT), but had limited impact on reducing global structural deviation (resulting in minimal RMSD improvement).
-Nevertheless, pLDDT alone is sufficient to demonstrate the high quality of the generated MSAs.
-4) Noise Issues: EvoDiff and MSAGPT produced MSAs that consistently underperformed compared to the original AF2 MSAs across all modes.
-This was particularly evident in low-homology protein scenarios, where the generated MSAs failed to accurately capture evolutionary distributions and often included irrelevant sequences.
-When concatenated with the original AF2 MSAs, these low-quality sequences introduced additional noise, disrupting the model’s inference.
-This aligns with the hypothesis presented in Section 3.4, where low-quality MSAs were predicted to introduce noise, ultimately impairing folding performance.
-5) Few/Zero-shot Consistency: In Few-shot scenarios, EvoDiff and MSAGPT performed relatively better due to the presence of initial MSAs that provided additional homology information to guide the generation process.
-In contrast, PLAME does not rely on initial MSA searches and directly generates high-quality MSAs through its evolutionary latent space.
-This enables PLAME to excel in Zero-shot scenarios while maintaining consistent performance in Few-shot cases.
-This advantage significantly reduces runtime costs while ensuring robust enhancements across all scenarios.
-Sequence quality assessment
+[…]
+[…]
+AF2 MSA 66.34 72.54 0.55 0.61 0.56 0.65 11.29 10.29 0.58 0.66 / /
+[…]
+4.1 Structure Benchmark Assessment
+We evaluated PLAME across three AF2 folding modes and AF3, using six structural metrics to assess MSA generation quality (See details in Table4).
+4.1.1 General Performance Comparison
+PLAME demonstrates consistent superiority across both zero-shot and few-shot scenarios against traditional MSA searching, AI-based searching, and AI-based generative methods, establishing a new paradigm for MSA generation without traditional homology search.
+In zero-shot settings, where proteins lack existing MSAs, PLAME achieves remarkable improvements with pLDDT scores reaching 71.50 in Mode3, significantly outperforming competing methods like EvoDiff (64.39) and MSAGPT (68.39).
+Moreover, the performance gap becomes even more pronounced in challenging scenarios: while EvoDiff and MSAGPT often introduce detrimental noise when their generated sequences are concatenated with original AF2 MSAs, PLAME consistently enhances folding quality.
+Interestingly, few-shot scenarios reveal that existing methods can partially recover performance when guided by initial homologous sequences, yet PLAME maintains its edge by generating more coherent evolutionary profiles that complement rather than interfere with existing MSAs.
+4.1.2 Mode-Dependent Performance Patterns
+The progression from Mode1 through AF3 reveals intriguing insights about the relationship between model sophistication and MSA augmentation benefits.
+Mode1 and Mode2 demonstrate the strongest relative improvements from PLAME-generated MSAs, with pLDDT gains of up to 5 points across different baseline methods.
+As configurations advance to Mode3 with structural templates, the enhancement effects become more nuanced—while absolute performance continues to improve, the marginal gains from MSA augmentation diminish because template information already captures substantial evolutionary constraints.
+This phenomenon reflects a fundamental trade-off in modern protein folding: as models become more powerful and incorporate diverse information sources, the additional value of synthetic MSAs decreases, though PLAME’s high-quality generations continue to provide meaningful contributions.
+The AF3 results further validate this trend, showing that PLAME maintains its effectiveness even with more advanced folding architectures, suggesting that high-quality virtual MSAs remain valuable complements to cutting-edge structural prediction methods.
+4.1.3 PLAME vs ESMFold: Bridging Efficiency and Accuracy
+The comparison with ESMFold reveals PLAME’s unique position in the protein folding landscape, offering a compelling alternative that combines computational efficiency with enhanced accuracy.
+While ESMFold achieves reasonable baseline performance (pLDDT of 66.26), PLAME progressively widens this gap as more sophisticated folding configurations are employed.
+In basic Mode1, PLAME shows modest improvements, but the advantage becomes substantial in Mode3 where PLAME reaches 71.50 pLDDT compared to ESMFold’s unchanged 66.26.
+This trend suggests that PLAME-generated MSAs provide increasingly valuable evolutionary context that more advanced folding models can effectively exploit.
+The consistent RMSD improvements across all modes further validate that PLAME’s virtual MSAs contribute meaningful structural constraints, enabling users to achieve AF2-level accuracy while maintaining the computational advantages of MSA-free approaches.
+4.2 Sequence quality assessment
+To evaluate generated MSA quality beyond structural perspectives, we conducted sequence-level analysis by fidelity and diversity metrics.
+This provides an additional critical gap—establishing criteria for understanding generated MSA quality.
+Figure 3 presents our comparative analysis.
+PLAME achieves superior evolutionary fidelity by closely mimicking the distributional characteristics of natural MSAs across all key metrics.
+The results reveal PLAME’s distributions align most closely with AF2 MSAs in Conservation Score, Gap Proportion, and Substitution Compatibility, demonstrating its ability to capture authentic evolutionary constraints.
 Figure 3: Comparison of sequence-based metrics for AF2 MSAs and MSAs generated by EvoDiff, MSAGPT, and PLAME.
-To evaluate generated MSA quality, we designed a set of metrics focusing on MSA fidelity and diversity.
-These metrics aim to investigate which factors correlate most strongly with protein folding performance, especially given the lack of standardized criteria for such evaluations.
-Our study provides an initial exploration into this problem.
-Figure 3 illustrates the evaluation results.
-PLAME-generated MSAs outperform other methods on fidelity metrics, with distributions closest to AF2 MSAs in Conservation Score, Gap Proportion, and Substitution Compatibility.
-Specifically, PLAME achieves higher Conservation Scores and Substitution Compatibility, reflecting its ability to better capture evolutionary and functional constraints.
-Moreover, its lower Gap Proportion indicates higher alignment completeness, which can be attributed to the latent evolutionary space providing richer homology information.
-These results highlight PLAME’s superior alignment with the evolutionary constraints of the target protein.
-We analyzed Alignment Entropy as a measure of diversity.
-While greater diversity is generally expected to enhance homologous information, our results show that PLAME’s diversity levels are closer to those of AF2 MSAs, rather than exceeding methods like EvoDiff.
-This supports our assumption in 3.4 that excessive diversity can introduce noise and diminish the information enrichment effect of generated MSAs.
-Therefore, achieving a balance between fidelity and moderate diversity is crucial for MSA generation models that struggle to capture high-quality distributions.
-4.2 PLAME as MSA-free ESMFold Adapter to AlphaFold2
-The comparative analysis between PLAME and ESM-Fold demonstrates PLAME’s growing advantage as the folding modes become more advanced.
-In Mode1 and Mode2, PLAME outperforms ESMFold in key metrics such as pLDDT and RMSD, with obvious gains in pLDDT.
-In Mode3, the introduction of structural templates amplifies PLAME’s advantage, achieving better results across all metrics.
-For instance, pLDDT of PLAME increases from 66.54 to 71.50, surpassing ESMFold’s 66.26.
-These results validate PLAME’s ability to merge ESMFold’s computational efficiency with AF2’s performance by generating high-quality virtual MSAs.
-Acting as an adaptive network, PLAME transforms ESMFold’s single-sequence inference into AF2’s MSA-based folding, integrating evolutionary information to enhance prediction accuracy.
-Thus, PLAME overcomes ESMFold’s limitations by introducing virtual MSA generation, achieving significant improvements in Mode3 and providing an efficient and accurate solution for protein structure prediction.
-Table 2: Ablation study over different MSA selection approaches.
-pLDDT GDT TMscore RMSD LDDT pTM
+This fidelity advantage manifests in higher Conservation Scores and Substitution Compatibility values, indicating that PLAME-generated sequences preserve functionally critical residues while incorporating biologically plausible substitutions.
+The significantly lower Gap Proportion validates PLAME’s approach, as the evolutionary latent space from ESM-2 provides richer homology information enabling more complete alignments.
+PLAME maintains diversity levels comparable to natural AF2 MSAs, supporting our hypothesis that excessive diversity introduces detrimental noise.
+Rather than maximizing diversity like EvoDiff, this measured approach aligns with our selection strategy principles, where balanced information enrichment proves more effective than naive sequence proliferation (Section 3.4).
+The findings suggest successful MSA generation requires maintaining the delicate balance between providing sufficient homologous information and avoiding noise from unconstrained sequence generation, positioning PLAME as a method that respects fundamental biological constraints.
+Table 2: Ablation study of HiFiAD on PLAME and other baselines.
+pLDDT  ($\uparrow$) GDT($\uparrow$) TMscore  ($\uparrow$) RMSD  ($\downarrow$) LDDT  ($\uparrow$) pTM  ($\uparrow$)
 Zero Few Zero Few Zero Few Zero Few Zero Few Zero Few
@@
 Top-down-Rec-16 63.43 63.10 0.52 0.52 0.51 0.57 11.97 12.15 0.55 0.58 0.46 0.49
-HiFiAD 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
-4.3 Ablation on Selection Methods
-To evaluate the impact of MSA quality (3.4), we perform an ablation study using different MSA selection strategies.
-The results are summarized in Table 2.
-The selection methods are based on sequence similarity metrics (Rec) and BLOSUM substitution scores (BLOSUM).
-Additionally, Top and Down refer to the highest- and lowest-scoring cases.
-The experiments lead to the following key observations.
-1) Effective HiFiAD filtering: HiFiAD achieves superior performance across all metrics compared to other filtering methods, demonstrating that the combination of high fidelity and balanced diversity is essential.
-2) High-fidelity: Compared to the Top-down-Rec method and random selection, HiFiAD identifies more high-fidelity cases, validating the importance of leveraging in-distribution MSAs.
-3) Diverse samples: Compared to BLOSUM-based methods and Top-Rec, HiFiAD selects more diverse samples.
-This balanced diversity prevents the conservation distribution from being overly deterministic, which could potentially correct the augmented MSA distribution.
-Among sequence-metric-based filtering methods, HiFiAD dynamically adjusts to Few- and Zero-shot scenarios, as well as varying MSA quality levels, making it a simple yet effective approach for MSA selection.
-More ablation results are shown in Table 9 in Appendix.
+EvoDiff-HiFiAD 58.24 60.89 0.46 0.49 0.46 0.54 13.74 12.39 0.51 0.56 / /
+MSAGPT-HiFiAD 60.16 62.63 0.48 0.52 0.48 0.57 12.54 12.18 0.53 0.59 / /
+DHR-HiFiAD 66.01 66.08 0.53 0.55 0.53 0.60 11.48 12.14 0.57 0.60 / /
+[…]
+4.3 Ablation Studies
+To validate our HiFiAD selection strategy, we conducted ablation experiments across different selection approaches and baseline methods.
+Table 2 compares various selection strategies and evaluates HiFiAD’s effectiveness on other generative methods.
+HiFiAD consistently outperforms alternative selection strategies by optimally balancing fidelity and diversity constraints.
+Compared to similarity-based methods (Top/Down-Rec) and substitution matrix approaches (BLOSUM-32), HiFiAD achieves superior performance with pLDDT scores of 66.54 in zero-shot settings, demonstrating the importance of jointly considering evolutionary fidelity and controlled diversity.
+The strategy effectively identifies high-fidelity sequences while maintaining sufficient diversity to prevent overly deterministic conservation patterns.
+HiFiAD automatically adapts to varying MSA quality levels and shot configurations, making it robust without requiring manual parameter tuning.
+When applied to competing baselines, HiFiAD consistently improves performance: EvoDiff benefits from a 58.24 to 60.89 pLDDT improvement, MSAGPT gains from 60.16 to 62.63, and DHR advances from 66.01 to 66.08.
+These improvements demonstrate that HiFiAD addresses fundamental challenges in MSA selection across all generative approaches.
+The consistent gains across different generation paradigms—from diffusion-based (EvoDiff) to autoregressive (MSAGPT) and retrieval-based (DHR) methods—validate that the fidelity-diversity trade-off represents a universal principle in MSA augmentation.
+The improvement margins correlate with baseline method quality, suggesting HiFiAD provides proportional benefits while maintaining relative performance hierarchy.
+Also, we conducted ablation study on MSA length (See Table3).
+PLAME shows overall improvement on all length ranges, where performs the largest improvement on 100-300 range.
+We believe this is because the MSA training data are mainly concentrated in this range (chen2024msagpt, ).
+Table 3: Ablation on protein length.
+Length Range pLDDT($\uparrow$) GDT($\uparrow$) TMscore($\uparrow$) RMSD($\downarrow$) LDDT($\uparrow$)
+AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
+AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
+AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
+PLAME <<100 74.12 0.63 0.52 7.49 0.61
+PLAME 100-300 65.55 0.53 0.58 11.58 0.58
+PLAME >>300 58.31 0.45 0.53 16.16 0.54
+Furthermore, we provide additional case studies on folding enhancement.
+More case studies on orphan de novo proteins (SectionE.5), protein failure cases (SectionE.4), selected protein cases with aligned structures (SectionE) can be found in the appendix.
+Figure 4: Case study of folding enhancement of PLAME, MSAGPT, and AF2 MSA on 8okh_B.
 5 Conclusion
-In this work, we introduced PLAME, the first model to leverage evolutionary embeddings for MSA sequence generation.
-Our approach bridges the gap between single-sequence inference and MSA-based methods, effectively enhancing protein folding performance.
-Our extensive evaluations demonstrate that PLAME-generated MSAs surpass existing methods in both conservation and diversity metrics, leading to significant improvements in structure prediction accuracy across various protein families.
-By functioning as both an MSA augmenter and an efficient AlphaFold adapter without the need for time-intensive MSA searches, PLAME establishes itself as a fast, accurate, and scalable solution for protein structure prediction.
-Furthermore, our proposed quality metrics and comprehensive experiments provide new insights into the relationship between MSA characteristics and folding performance, advancing our understanding of how sequence information translates to structural accuracy.
+In this study, we introduce PLAME, the first model to leverage evolutionary embeddings for MSA generation and apply it to protein folding enhancement.
+Our approach bridges the gap between single-sequence inference and MSA-based methods, effectively improving protein folding performance.
+Evaluation results demonstrate that PLAME-generated MSAs outperform existing methods in both conservation and diversity metrics, significantly enhancing structural prediction accuracy across different protein families.
+PLAME serves as both an MSA enhancer and an efficient AlphaFold adapter without requiring time-consuming MSA searches, providing a fast, accurate, and scalable protein structure prediction solution.
+Additionally, our proposed quality metrics and experiments offer new insights into the relationship between MSA features and folding performance.
 Appendix A Proof of Theorem
@@
 Sampling was executed on an A40 GPU.
-Appendix C Extensive Benchmark Results
-C.1 Comparison to AI-based MSA Retrieval Methods
-We further compare the folding enhancement by MSAs retrieved by AI-based approaches [19].
-Without filtering, DHR’s performance is generally higher than AF2 MSA, EvoDiff, and MSAGPT, demonstrating the effectiveness of applying evolutionary embeddings like ESM2.
-Zero/Few pLDDT GDT TMscore RMSD LDDT
-DHR Zero 63.64 0.51 0.52 12.04 0.55
-DHR Few 62.60 0.52 0.57 11.92 0.59
-PLAME Zero 66.54 0.53 0.53 11.48 0.57
-PLAME Few 66.08 0.54 0.58 12.14 0.60
-Table 3: Comparison of folding enhancement to AI-based MSA retrieval method.
-C.2 Comparison on Inference Speed and Memory Usage
+Appendix C Related Works
+Protein Structure Prediction
+Protein structure prediction methods fall into three main categories: physics-based, homology-based, and deep learning approaches.
+Physics-based methods, such as AMBER and CHARMM, use molecular physics and energy optimization to simulate protein folding [10, 20].
+While offering detailed folding insights, they are computationally expensive and sensitive to initial conditions, often yielding suboptimal results [21, 22, 23].
+Homology modeling tools, like Rosetta and HHpred, use MSAs and evolutionary data to predict structures by refining templates from known experimental structures [9, 11].
+These methods perform well with suitable templates but struggle with orphan proteins and low-homology families [8, 1].
+Deep learning-based methods, such as AlphaFold2 and OmegaFold, use advanced neural architectures and protein templates to achieve near-experimental accuracy with greater speed and scalability [3, 5, 24].
+Despite their success, they still depend on high-quality MSAs and struggle with low-homology proteins.
+AlphaFold-based Enhancement
+Building on AlphaFold’s success, researchers have developed methods to refine specific modules, aiming to improve accuracy or efficiency.
+These advancements can be grouped into three main categories.
+The first category focuses on homology expansion techniques, such as MMSeq2 and DeepMSA2, which expand the evolutionary search space to enhance prediction accuracy.
+However, these methods often slow down inference despite their modest performance gains [25, 26, 27, 28].
+The second category targets search acceleration, with methods like ColabFold and ESMFold bypassing the MSA search process to enhance computational efficiency.
+However, this speedup often results in incomplete evolutionary data, potentially reducing prediction accuracy [6, 29].
+The third category leverages generative models to capture protein homology and augment input data, especially for orphan proteins and low-homology families.
+While promising in specific scenarios, these models struggle with extremely limited evolutionary signals, and their artificial sequences often deviate from traditional MSA distributions, limiting broader applicability [16, 14, 13, 15].
+Appendix D Comparison on Inference Speed and Memory Usage
 To further demonstrate PLAME’s efficiency, we calculated the inference time and memory cost of each method.
@@
 Table 4: Comparison on inference speed and memory.
-C.3 Ablation on HiFiAD Filtering
-To demonstrate the effectiveness of HiFiAD, we conducted ablation experiments on filtering methods for other baselines.
-Among the benchmarks, PLAME maintains SOTA performance, and DHR-HiFiAD and PLAME show comparable performance across multiple metrics.
-Several metrics of EvoDiff, MSAGPT, and DHR (AI-based MSA retrieval methods) all demonstrate improvement after HiFiAD filtering compared to their original filtering methods, which directly proves the benefit of HiFiAD rules for MSA in AF-series structure prediction.
-This insight can help future work develop higher-quality MSA sequences.
-Zero/Few pLDDT GDT TMscore RMSD LDDT
-DHR-HiFiAD Zero 66.01 0.53 0.53 11.48 0.57
-DHR-HiFiAD Few 66.08 0.55 0.60 12.14 0.60
-EvoDiff-HifiAD Zero 58.24 0.46 0.46 13.74 0.51
-EvoDiff-HifiAD Few 60.89 0.49 0.54 12.39 0.56
-MSAGPT-HifiAD Zero 60.16 0.48 0.48 12.54 0.53
-MSAGPT-HifiAD Few 62.63 0.52 0.57 12.18 0.59
-Table 5: Ablation on HiFiAD filtering.
-C.4 Ablation on Protein Length
-We listed the performance of PLAME and AF2 MSA in different length ranges, and found that PLAME shows overall improvement, with the largest improvement in the 100-300 range, followed by the >>300 and <<100 ranges.
-We believe this is due to our MSA training data being mainly concentrated in the 100-300 length range.
-Length Range pLDDT GDT TMscore RMSD LDDT
-AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
-AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
-AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
-PLAME <<100 74.12 0.63 0.52 7.49 0.61
-PLAME 100-300 65.55 0.53 0.58 11.58 0.58
-PLAME >>300 58.31 0.45 0.53 16.16 0.54
-Table 6: Ablation on protein length.
-Appendix D Extensive Case Studies
-D.1 Case Study on Successful Designs
-To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix G.
+Appendix E Extensive Case Studies
+E.1 Case Study on Successful Designs
+To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix H.
 We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.
@@
 These represent an area requiring further improvement.
-D.2 AlphaFold3 Folding Experiment
-To further validate PLAME is an universal MSA generation framework for different folding software, we conduct extensive experiment on AlphaFold3 [5].
-We use the same generated MSAs filtered by HiFiAD, augmenting the original MSAs for folding enhancement.
-pLDDT GDT TMscore RMSD LDDT pTM
-AF2 MSA 68.872 0.578 0.596 10.818 0.617 0.529
-PLAME 70.887 0.578 0.595 10.740 0.623 0.539
-Table 7: Comparison of folding enhancement based on AlphaFold3
-Results follow the same trend as AF2 Mode 3.
-The stronger the folding baseline, the smaller the performance gain.
-Improvements are mainly in pLDDT and RMSD, indicating our enhanced MSA primarily improves local structural regions while preserving the global protein architecture.
-D.3 Folding Enhancement on Average Proteins
+E.2 Folding Enhancement on Average Proteins
 To probe the effectiveness of PLAME on average proteins, we firstly build a dataset from PDB validation set with 36 proteins.
@@
 The relatively small improvements observed when applying our method to the average protein MSA can be attributed to the fact that these original MSAs already provide sufficient evolutionary information to AlphaFold2’s MSA Transformer, thus limiting the potential impact of additional augmentation.
-D.4 Further Ablation on MSA Filtering
+E.3 Further Ablation on MSA Filtering
 We further validate the effectiveness of filtered high-quality MSAs by comparing the performance with the more randomly selected MSAs (64 for each protein).
@@
 From the results, more co-evolutionary information may lead to better local geometric conformation, but it will disturb the modeling of the global conformations due to the bias during generation.
-D.5 Failure Case Analysis
+E.4 Failure Case Analysis
 Other than analyzing successful cases, we analyzed four representative failure cases (3bog_B, 7sxb_A, 8gzu_AN, 8gzu_T3) with the largest performance drops, which includes three zero-shot and one few-shot examples.
@@
 Table 7: Comparison of folding enhancement on failure cases.
-Among the visualized MSA cases, we observed that generated MSAs contained extremely similar sequences (>90% similarity).
+Among the visualized MSA cases, we observed that generated MSAs contained extremely similar sequences (>>90% similarity).
 Specifically, these high-similarity sequences caused all sites to appear more conserved, resulting in a lack of covariation patterns necessary for AlphaFold2 to infer structural contacts.
@@
 Notably, for 3bog_B and 8gzu_T3, the generated high-similarity MSAs further enhanced the conservation of already conserved regions, which consequently led to improvements in global metrics.
-D.6 De novo Protein Folding Enhancement
+E.5 De novo Protein Folding Enhancement
 We conduct further experiments on De Novo protein cases, where almost of them are orphan.
-Examples of de novo proteins include 8SK7 (RFDiffusion), 8TNM/8TNO (Chroma), and 8CYK (ProteinMPNN).
+Examples of de novo proteins include 8SK7 (RFDiffusion [38]), 8TNM/8TNO (Chroma [39]), and 8CYK (ProteinMPNN [40]).
 We followed the same augmentation pattern as the main experiment.
@@
 From Table 8, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics.
-The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $70\%$ similarity to the query sequences.
+The generated MSA visualizations in Figures 5 and 6 reveal that most generated sequences maintain $70\%$ similarity to the query sequences.
 This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset.
@@
 Figure 6: Comparison of structure enhancement of De Novo proteins.
-Appendix E Discussion
-E.1 Limitations
+Appendix F Discussion
+F.1 Limitations
 Recent advancements in MSA generation models have shown promising results in enhancing protein folding predictions.
@@
 Developing direct and robust quality assessment metrics will be crucial for systematically evaluating and improving MSA generation methods, enabling the selection of high-quality MSAs for specific applications and paving the way for next-generation models with enhanced accuracy, broader applicability, and greater biological relevance.
-E.2 Social Impact
-PLAME advances multiple sequence alignment (MSA) generation for proteins, offering significant potential benefits for protein structure prediction and engineering applications.
-This capability could accelerate therapeutic protein development, enhance drug design processes, and facilitate the exploration of novel functional protein sequences.
-However, we acknowledge important limitations and risks associated with this technology.
-The model may occasionally generate hallucinated sequences that could mislead downstream protein design efforts if not properly validated.
-Furthermore, like many powerful biotechnology tools, there exists potential for misuse in designing harmful biological entities.
-Appendix F Structure Comparison Visualization
+Appendix G Structure Comparison Visualization
 Figure 7: Structure comparison visualization of 8ehb_F.
-Figure 7: Structure comparison visualization of 8okh_B.
 Figure 8: Structure comparison visualization of 8b4k_C.
@@
 Figure 11: Structure comparison visualization of 8okw_B.
-Appendix G Augmented MSA Visualization
+Appendix H Augmented MSA Visualization
 To provide an intuitive understanding of the MSAs generated by PLAME, we selected several representative cases for visualization.
@@
 Figure 18: Augmented MSA visualization of 8e0n_F.
-Appendix H Failure Case MSA Visualization
+Appendix I Failure Case MSA Visualization
 Figure 19: Failure Case MSA visualization of 7sxb_A.